### v1

# MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning

Abstract
Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities beyond the pattern-matching strengths of LLMs. LLMs struggle with planning tasks because of their reliance on associative reasoning, inability to self-verify, and inconsistent constraint awareness. We propose Multi-Agent Collaborative Intelligence (MACI), a framework centered on a meta-planner (MP) that orchestrates multiple agents to generate planner templates that define roles and constraints. These planners produce actionable workflows of role nodes and dependency constraints, enabling advanced temporal reasoning and adaptability.
MACI’s three-tier architecture includes a meta-planning module for planner construction, common agents for general reasoning, and specialized agents for domain expertise. By decoupling planning from validation, it overcomes key LLM limitations. Evaluations demonstrate MACI’s effective constraint satisfaction, conflict detection, and reasoning, positioning it as a robust solution for complex reasoning and planning tasks.
Machine Learning, ICML
1 Introduction
The development of artificial general intelligence requires a shift from pattern matching to deliberate analytical reasoning. Large Language Models (LLMs) excel at pattern matching but lack deliberative reasoning required for planning tasks (Kahneman, 2011). These capabilities become crucial for tasks requiring sustained attention, awareness of constraints, and reasoning about future states.
1.1 Limitations of LLMs in Planning
Our evaluation reveals three critical limitations in current LLMs that prevent effective planning.
Metacognitive Limitations in Self-verification
LLM limitations in self-verification extend beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967). Though formal systems cannot prove their own consistency within their framework, LLMs face greater challenges due to their probabilistic nature and lack of logical foundations (detailed arguments in Section 2). This limitation provides theoretical justification for distributed validation, where specialized agents perform independent verification of planning processes.
Attention-based Processing Bias
In complex planning scenarios, LLMs demonstrate a form of cognitive tunneling where recent context dominates earlier-established constraints. For example, when planning a multi-stop journey, an LLM might focus on optimizing the final leg’s arrival time while losing sight of earlier constraints about vehicle availability or required rest periods. This attention bias leads to locally optimized but globally invalid solutions, undermining the overall feasibility of the plan. The inability to maintain consistent awareness of all constraints throughout the planning process represents a fundamental limitation in current architectures.
Common Sense Knowledge Integration
LLMs often overlook practical constraints that humans naturally consider. In travel planning, they might ignore factors such as airport security wait time or luggage claim delays. These omissions are the result of an inadequate integration of common sense knowledge into the planning process. Although humans automatically incorporate such practical considerations, LLMs require explicit specification of these constraints, often leading to incomplete or impractical plans.
Table 1: Comparison of Validation Limitations and Solutions
Gödel’s Incomp. Theorems LLM Validation Limitations MACI’s Solution
A formal system cannot prove its own consistency LLMs cannot reliably validate their own outputs due to probabilistic foundations MACI uses external validation agents to bypass self-referential loops
Consistency proofs require a “higher” system Validation requires external frameworks (e.g., logic engines, human oversight) Distributed validation acts as the “higher" system for LLM outputs
Self-reference leads to paradox or incompleteness LLMs hallucinate or contradict themselves when critiquing their own outputs Meta-planning separates planning and validation into distinct agent roles
1.2 The MACI Framework
To overcome the inherent limitations of single-model approaches in planning tasks, we introduce the Multi-Agent Collaborative Intelligence (MACI) framework. At the core of MACI lies the meta-planner (MP), a high-level orchestration mechanism that collaborates with multiple agents to construct a specialized planner for temporal reasoning and scheduling. By enabling agents to work collectively at the meta-level, MACI ensures robust and adaptive planning across complex domains.
Meta-Planning Module
The meta-planning module serves as the cornerstone of MACI, analyzing task requirements and dynamically generating a planner specifically tailored to address complex challenges. This planner, in turn, produces a workflow composed of nodes representing roles (e.g., cook, driver, supervisor) and dependencies representing constraints (e.g., temporal, spatial, resource). Unlike traditional multi-agent systems, MACI’s meta-planner orchestrates a dynamic network of agents, validates their interactions, and enforces protocols to ensure comprehensive planning and execution. This layered approach—planning to generate a planner that produces actionable workflows—positions MACI as a forward-looking solution to real-world planning problems.
Common Agents
Common agents form the foundation of MACI’s general-purpose planning capabilities. They address critical tasks such as practical reasoning, constraint validation, and robustness analysis. For example, the Common Sense Integration agent identifies implicit constraints, the Constraint Validation agent ensures solution feasibility, and the Plan Evaluation agent measures performance metrics. These agents collectively maintain both explicit and implicit constraints while ensuring compliance and safety across domains.
Specialized Agents
Specialized agents extend MACI’s capabilities by addressing domain-specific requirements. Instead of overburdening a single planner, MACI deploys focused agents for tasks such as travel scheduling, financial risk assessment, or logistics optimization. These agents integrate domain expertise with the foundational work of common agents, ensuring precise and effective solutions for specialized planning challenges.
1.3 Contributions
This work introduces MACI as a novel approach to collaborative planning. Preliminary evaluations in travel planning scenarios demonstrate MACI’s superiority in constraint satisfaction, conflict detection, and practical reasoning compared to single-LLM approaches. Key contributions of this paper include:
  1. 1.
    Identifying limitations of single-LLM approaches through systematic experimentation.
  2. 2.
    Developing a three-tier architecture for distributed planning and validation.
  3. 3.
    Designing MACI with common and specialized agents for constraint management and reasoning.
  4. 4.
    Demonstrating MACI’s effectiveness in travel planning scenarios, achieving significant improvements over single-LLM systems.
The remainder of this paper is organized into five sections. Section 2 reviews related MAS frameworks and explains why they fall short for the task at hand. Section 3 examines the limitations of LLMs through case studies. Section 4 details the specifications of the MACI framework. Section 5 evaluates MACI’s performance in comparison to state-of-the-art LLMs. Finally, Section 6 discusses implications for future research and alternative perspectives.
2 Related Work
The development of MACI builds on theoretical insights from formal systems and addresses limitations of current multi-agent architectures. Gödel’s second incompleteness theorem (Kennedy, 2008; Gödel, 1967) established that no consistent formal system can prove its own consistency. We extend this insight to LLMs, which operate on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation. Table 1 illustrates how MACI’s distributed validation architecture addresses these fundamental limitations.
Current multi-agent systems (MAS) primarily serve as integration platforms for coordinating multiple LLMs. Notable frameworks include Microsoft’s AutoGen (Wu et al., 2024), the Multi-LLM Agent Debate Framework (Du et al., 2023; Chang, 2023, 2024a, 2024b), LangGraph and CrewAI (LangChain AI, 2024; Moura, 2024), XAgent (Xia et al., 2023), and CAMEL (Li et al., 2023). Although these frameworks excel at agent coordination, they focus on task distribution rather than comprehensive constraint management required for complex planning.
MACI’s meta-planning module (Section 4.1) addresses these limitations by acting as a metasystem that coordinates specialized agents. By decoupling planning and validation functions, MACI creates a validation framework analogous to the higher-order system required to demonstrate formal consistency while maintaining awareness of interdependent constraints.
3 Case Study: Can LLMs Handle Planning
This section begins by presenting various types of planning, including strategic approaches and tactical methodologies. We then present experimental results on a scheduling problem to illustrate how current LLMs fall short. Based on our diagnoses, we propose remedies in Section 4.
Planning Types
Planning can be classified into seven types. The first three are considered more strategic and the rest are tactic methodologies.
  1. 1.
    Sequential planning: Creates time-ordered schedules meeting goals/constraints via formal temporal representations (Allen & Hayes, 1989)
  2. 2.
    Proactive planning: Anticipates future scenarios to preemptively address issues through deliberate foresight (Cox & Veloso, 1998)
  3. 3.
    Retrospective planning: Improves future plans by analyzing past experiences via case-based reasoning (Kolodner, 1993)
  4. 4.
    Adaptive planning: Adjusts plans to changing conditions while preserving core constraints (Hammond, 1990)
  5. 5.
    Reactive planning: Prioritizes immediate actions over long-term goals in dynamic, uncertain environments (Georgeff & Lansky, 1987)
  6. 6.
    Predictive planning: Uses data-driven models to forecast scenarios and optimize decisions (Kushmerick et al., 1995)
Problem Statement
We conduct experiments using a Thanksgiving dinner planning problem designed as follows:
Initial Setup:
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 1:00 PM Eastern time.
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James needs to rent a car after landing
  • Emily doesn’t drive, needs pickup from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven
  • Side dishes require 2 hours of preparation
  • Travel time between home and BOS airport is one hour
  • Travel between BOS and grandma home is one hour
  • Travel between home and grandma home 30 minutes
Planning Question Set:
  1. 1.
    At what time should cooking start?
  2. 2.
    Who should pick up Emily and when?
  3. 3.
    When should Grandma be picked up by who?
Using this problem, three key types of planning conducted by GPT4o and DeepSeek are examined: sequential, proactive/reactive, and retrospective.
Table 2: DeepSeek Proposed Schedule
Time Task Assigned
12:00pm Sarah starts preparing side dishes. Sarah
1:00pm James arrives at the airport, begins car rental (takes 30-60 minutes). James
2:00pm James is ready to pick up Emily; Sarah and James put the turkey in the oven. J, S
2:30pm Emily lands at BOS. Emily
3:00pm James arrives at airport to pick up Emily J
3:00pm Michael arrives at home and leaves to pick up Grandma. Michael
3:15pm Michael at Grandma’s home. M
3:45pm Grandma arrives home; Michael returns. M
4:00pm James and Emily arrive home. J, E
6:00pm Dinner is served. All
3.1 Study #1 Sequential Planning
GPT-4o and DeepSeek struggled with real-world travel logistics, omitting key human considerations such as luggage claim time, rental car processes, and spatial relationships between airport terminals and the rental center (Table 2). These gaps forced manual constraint additions, highlighting LLM’s inability to integrate experiential knowledge, a gap addressed by the MACI common sense agent.
DeepSeek’s schedule further revealed spatial-temporal errors: 1) Spatial: Assumed James drove home immediately after renting a car at Boston Logan, ignoring his airport location while awaiting Emily; and 2) Temporal: Directed Michael to return home before heading to Grandma’s, missing the optimal direct route from NYC.
Table 3: GPT4o Proposed Schedule
Time Task Assigned
1:00pm James lands in Boston James
2:00pm Turkey goes into the oven Sarah
2:00pm James finishes car rental J
2:30pm Emily lands at BOS Emily
2:30pm James picks up Emily at airport J
3:00pm Michael arrives home Michael
4:00pm Side dishes preparation starts S, M
5:00pm Michael leaves to pick up Grandma M
5:30pm Michael arrives home with Grandma M
6:00pm Dinner is served All
Table 3 shows the GPT-4o schedule, which appears feasible but contains two critical errors in the case of adaptive planning required for emergency: 1) Arithmetic: Incorrectly calculates Grandma’s round-trip driving time as 30 minutes (vs. 30×2 minutes); and 2) Over-Constraint: Assumes only Sarah must watch the oven (vs. “someone”), creating brittleness under reduced slack time (e.g., delays).
Analysis (Appendix A) links both errors to flawed reasoning in constraint interpretation.
Diagnoses: Common-Sense Constraints and Isolated Processing Syndrome
Current LLM systems require explicit specification of real-world constraints that humans consider common sense, highlighting a limitation in their planning capabilities. Furthermore, we identified what we term isolated processing syndrome, where LLMs handle sub-tasks independently without maintaining awareness of overall constraints. This syndrome manifests itself in two critical ways: the system either misses obvious optimizations or proposes solutions that violate the stated constraints, leading to an infeasible or suboptimal plan.
3.2 Study #2 Proactive and Reactive Planning
Real-world scenarios do not always follow plans precisely. Robust systems require contingency planning for factors such as weather, traffic, or airline changes. These cascade through schedules, demanding adaptive replanning.
Table 4: GPT4o Revised Thanksgiving Schedule. Hazard! No one home watch oven between 3:00pm and 4:00pm.
Time Task Assigned
2:00pm Turkey placed in oven (4-hour cooking time begins) Sarah
3:00pm Michael arrives home Michael
Michael departs to pick up Emily from airport Michael
3:00pm Sarah departs to pick up Grandma Sarah
3:30pm Arrive at Grandma’s house Sarah
4:00pm Arrive at airport for Emily’s pickup Michael
Sarah home with Grandma -
James’s flight lands James
Begin side dish preparation Sarah
4:30pm James completes car rental process James
5:00pm Michael returns home with Emily -
5:30pm James arrive home -
6:00pm Thanksgiving dinner served Everyone
To evaluate LLMs’ dynamic replanning, we introduce a major disruption in our Thanksgiving scenario: James’s flight is delayed by 3 hours (arrival 4:00 PM vs. 1:00 PM). This forces adjustments to pickups, meal prep, and coordination while preserving original constraints.
LLM responses reveal critical flaws: 1) DeepSeek violates core constraints by unjustifiably delaying dinner to 7:00 PM (vs. the 6:00 PM deadline); and 2) GPT-4o (Table 4) commits a safety violation: leaving the oven unattended, despite explicit constraints. These errors highlight LLMs’ inability to reliably maintain and validate constraints during replanning, even with full information.
Diagnosis: Attention Narrowing
Claude detects constraint violations in other LLMs’ plans but both GPTo4 and DeepSeek struggle with self-validation, revealing an asymmetry in error detection. LLMs often embed flawed interpretations of constraints during planning (e.g., rigidly interpreting “someone must be in the house” when the turkey is in the oven), propagating errors through their frameworks.
Two key limitations emerge: 1) Attention narrowing: Overfocus on objectives (e.g., arrival times) causes neglect of critical constraints (e.g., fire safety); and 2) Solution rigidity: Once a constraint is satisfied (e.g., assigning Sarah to oven duty), LLMs treat it as a fixed context, failing to explore alternatives.
More specifically, GPT-4o assigned Sarah to monitor the oven, but missed reallocating this task to Grandma earlier, preventing Sarah from serving as an additional driver, a missed efficiency gain.
3.3 Study #3 Retrospective Planning
Retrospective planning improves future strategies by learning from past decisions. Both DeepSeek and GPT-4o lacked flexibility even after error correction: DeepSeek relied on rigid templates, and GPT-4o’s initial assignment (Sarah homebound in Table 3) limited later options.
Missed Opportunities During James’ 3-Hour Delay:
  • Sarah could have picked up Grandma earlier, freeing her for contingencies like Emily’s transport.
  • Michael could have left NYC at 10:00 AM to reach Emily by 2:30 PM.
In Section 4, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
3.4 Summary of LLM Limitations in Planning
Our analysis reveals three core limitations in current LLMs and reasoning methods (CoT (Wei et al., 2022), ToT (Yao et al., 2023)):
Metacognitive Limitations
LLMs struggle with self-validation and constraint awareness. While external LLMs detect errors in others’ plans, planners consistently miss their own violations (e.g., GPT-4o rigidly assigned Sarah to oven duty without considering Grandma’s availability). Key causes are:
  1. 1.
    Pattern-matching optimization vs analytical validation
  2. 2.
    No belief-state tracking during reasoning
  3. 3.
    Single-solution focus vs comparative analysis
Current reasoning methods exacerbate these issues by operating within the same flawed cognitive framework.
Attention Bias
Transformer architectures prioritize recent context, creating: 1) Narrowing: Recent constraints (arrival times) overshadow earlier ones (oven safety); and 2) Isolated Processing: Sub-tasks addressed without holistic awareness
Common Sense Gaps
LLMs miss implicit real-world knowledge (luggage claim times, rental car logistics) requiring explicit specification of human-obvious constraints (airport-terminal proximity).
4 MACI Framework Specification
MACI implements a three-layer architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning. The top two layers represent key innovations, while the bottom layer leverages existing multi-agent system (MAS) infrastructure.
4.1 Three-Layer Architecture
Meta-Planner: The Planning Generator
The meta-planner $\mathcal{MP}$ functions as a higher-order planner that generates task-specific planning systems:
$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},$
where $\mathbf{W}$ represents a planning system composed of specialized, coordinated agents. Similar to how a compiler generator produces compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements. It analyzes objectives, identifies required roles and dependencies, selects appropriate agents, and establishes interaction protocols. This produces a complete workflow template defining both the planning state space and the agent coordination mechanisms needed to solve the given task.
Agent Repository Layer
This layer maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface. By dividing cognitive tasks among agents, this architecture ensures a complete knowledge of the constraints without overwhelming individual components.
System Infrastructure Layer
Built on open-source MAS frameworks, the infrastructure layer supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling.
4.2 Agent Repository Layer Design
4.2.1 Agent Registration and Search Protocol
Agents register in the repository with five attributes: input/output protocol ($P$), type ($t$), capability vector ($c$), context window size ($w\leq 1k$ tokens) and user rating ($r$). These attributes enable $\mathcal{MP}$ to match task requirements to agent capabilities efficiently. $\mathcal{MP}$ searches for agents by: 1) matching task requirements to capability descriptions ($c$), 2) filtering by agent type ($t$) and user ratings ($r$), and 3) verifying protocol compatibility ($P$).
4.2.2 State Space and Agent Design
General State Space
Table 5: State Space Dimensions and Components
Dimension Core Components Example States
Who
• Actor ID
• Current Roles
• Role Constraints
• Role Changes
• Driver vs Passenger
• Supervisor vs Worker
• Buyer vs Seller
• Multiple role conflicts
Where
• Current Position
• Target Location
• Transition Points
• Access Rules
• Physical locations
• Virtual positions
• State transitions
• Boundary constraints
When
• Time Points
• Duration
• Deadlines
• Probability
• Event timestamps
• Process duration
• Completion times
• Delay likelihood
What
• Methods
• Requirements
• Constraints
• Costs
• Tools/Vehicles
• Tickets/Permits
• Capacity limits
• Time-money trade-offs
Why
• Rationale
• Dependencies
• Risks
• Alternatives
• Decision basis
• Causal chains
• Failure modes
• Backup plans
Table 5 presents five dimensions of state space, Who, Where, When, What, and Why, that form the basis of planning tasks. These dimensions are general enough to apply across domains, with travel planning serving as an illustrative example.
Common Agents
Common agents are designed to handle core planning and validation tasks that are universally applicable across domains. These agents provide foundational capabilities that ensure the consistency, feasibility, and robustness of planning systems, regardless of the specific application. Their roles align with the general state space dimensions of Who, Where, When, What, and Why, ensuring that all essential aspects of planning are addressed. For a list of common agents, please refer to Appendix D.2.
These agents work collaboratively to ensure robust planning and validation across diverse domains. Although the list of common agents is not exhaustive, it provides a strong foundation to address most planning tasks. New common agents can be introduced as the need arises, provided they offer capabilities that can be reused across multiple domains. For example, an Adaptation Agent could be added to dynamically adjust constraints in response to evolving task environments.
Specialized Agents
Specialized agents extend the capabilities of MACI by addressing domain-specific requirements that are not generalizable across all tasks. These agents are designed to implement detailed knowledge and expertise tailored to specific applications, ensuring that MACI remains adaptable and effective for diverse planning challenges. Specialized agents typically focus on tasks requiring unique constraint handling, optimization, or analysis.
Examples of specialized agents include the following.
  • Travel Planning Agent: Optimizes itineraries, manages transportation options, and ensures seamless transitions between locations.
  • Financial Risk Agent: Monitors market trends, assesses risks, and validates investment strategies.
  • Healthcare Resource Agent: Allocates medical supplies, schedules staff, and optimizes patient care pathways.
  • Supply Chain Agent: Manages logistics, tracks inventory, and optimizes distribution routes.
Specialized agents adhere to architectural principles that maintain consistency and integration within MACI:
  • Restricted Context Windows: Each agent operates within a limited context to ensure efficiency and focus ($w\leq 1k$ tokens).
  • Protocol Buffer Interfaces: Standardized interfaces enable seamless communication between agents.
  • Coordination with Common Agents: Specialized agents leverage the capabilities of common agents for validation, compliance, and constraint management.
  • Domain-Specific Constraints: Specialized agents are equipped to handle unique requirements and challenges specific to their domain.
4.2.3 Agent Coordination and State & Constraint Management
MACI ensures seamless coordination between agents across state dimensions. For example, in travel planning, when a Schedule Agent detects a flight connection constraint, it triggers updates to Location, Resource, and Role Manager Agents. These updates propagate through a state transition protocol, where the proposed state change is validated against global and local constraints:
$T:S_{t}\rightarrow S_{t+1}\text{ subject to }C_{\text{global}}\land C_{\text{local}}$
4.2.4 Validation Protocols
The validation agent ensures state consistency by decomposing global constraints into dimension-specific checks. For example, when booking a flight, the temporal agent validates the temporal constraint, the spatial feasibility is checked by the location agent, and the availability of seats is validated by the resource agent. Common sense agents identify overlooked constraints, such as luggage claiming time, inter-terminal walking time, and rental car pickup time. If validation fails, the system triggers a recovery process to generate alternative solutions. Validation protocols can be updated when unexpected events occur to improve retrospective planning.
4.2.5 Other Considerations
Due to space limitation, we document in Appendix C additional considerations including cross-domain generalization, enhancement and integration of common sense knowledge, scalability, and a list of future enhancements when more applications will be supported.
4.3 Meta-Planner: Planning a Planner to Plan
The meta-planner $\mathcal{MP}$ operates as a higher-level planner that constructs other planners, which are actionable workflows for specific tasks.
4.3.1 The Meta-Planner Algorithm
$\mathcal{MP}$ is specified in Algorithm #1. A list of all symbols used in the specifications is provided in Table 20 in Appendix D.
A workflow network $\mathbf{W}=(\mathcal{N},\mathcal{E})$, represented as a directed graph, consists of roles as nodes and their dependencies as edges. The nodes $n\in\mathcal{N}$ denote the roles performed by human actors (e.g. driver, cook, supervisor), while the edges $e_{ij}\in\mathcal{E}$ between the nodes $n_{i}$ and $n_{j}$ capture the monitored dependencies (e.g., temporal, spatial constraints, data dependencies and safety requirements for supervision).
$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n}\text{and }\mathcal{E}=A^{*}_{e}.\vspace{-.1in}$ (1)
Algorithm 1 $\mathcal{MP}$: Planner for Planning a Plan
0:  Objectives $\mathcal{O}$, explicit constraints $\mathcal{C_{E}}$, agent pool $\mathbf{A}$, people $\mathcal{P}$, metrics $\mathcal{M}$
0:  Optimized workflow $\mathbf{W^{*}}=(\mathcal{N},\mathcal{E})$ (Eq. 1)
  // Phase 1: Network Construction
  1. $\mathcal{N}$: Extract roles & capabilities from $\mathcal{O}$ (Eq. 2)
  2. $\mathcal{E}$: Identify node dependencies from $\mathcal{C_{E}}$ (Eq. 3)
  // Phase 2: Agent Assignment
  3. Assign role agents: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
  4. Assign edge agents: $\forall e_{ij}\in\mathcal{E}$, $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
  // Phase 3: Iterative Refinement
  while improvement in $V(\mathbf{W},\mathcal{M})$ do
     // Role Assignment Optimization
     for all $n\in\mathcal{N}$ do
        Update role-person mappings via $f_{\text{role}}(n,\mathcal{P})$
     end for
     // Dependency Monitoring
     for all $e\in\mathcal{E}$ do
        Verify dependencies via assigned edge agents
     end for
     // Validation
     if $V(\mathbf{W_{new}},\mathcal{M})>V(\mathbf{W_{current}},\mathcal{M})$ then
        $\mathbf{W_{current}}\leftarrow\mathbf{W_{new}}$
     end if
  end while
  return $\mathbf{W^{*}}=\mathbf{W{current}}$
$\mathcal{MP}$ constructs a planner in three phases. First, it builds the network structure by identifying roles and their dependencies, forming the foundation of the workflow.
Second, it assigns appropriate agents to manage both roles and dependencies. Each node $n\in\mathcal{N}$ is managed by a designated role agent $\alpha_{n}$. Similarly, each edge $e_{ij}\in\mathcal{E}$, representing a dependency between nodes, is managed by an agent $\alpha_{ij}$, which connects the node agents $\alpha_{i}$ and $\alpha_{j}$. These edges are enriched with metadata that may include details such as data types, communication protocols, timing constraints, and security requirements.
Finally, the generated planner iteratively refines the workflow. This process involves optimizing role assignments and continuously monitoring dependencies until the validation score stabilizes, ensuring a well-validated workflow.
4.3.2 Meta-Planning Components
Let us clearly distinguish between roles (requiring human qualifications) and monitoring tasks (requiring capabilities).
Role and Qualification Analysis
$\mathcal{MP}$ analyzes objectives $\mathcal{O}$ to identify required roles and their qualifications:
$\text{map}_{\text{role}}:\mathcal{O}\rightarrow\{(n_{i},q_{i})\}$ (2)
where $n_{i}$ represents a role and $q_{i}$ its required qualifications. For example, $(n_{\text{driver}},q_{\text{license}})$ or $(n_{\text{cook}},q_{\text{experience}})$.
Constraint Management
The framework maintains a global constraint set $C$ composed of:
$C=C_{E}\cup C_{I}\cup C_{D}$ (3)
where $C_{E}$ represents explicit constraints from problem specification, $C_{I}$ denotes implicit constraints identified by Common Sense Agent, and $C_{D}$ represents derived constraints from agent interactions.
Agent Assignment
Two types of agents, node and edge, are selected from the repository $\mathbf{A}$:
1. Node Monitoring Agents:
$A^{*}_{n}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{n_{j}}\text{dist}(q_{j},A_{i}.\text{capabilities})$ (4)
These agents verify role qualifications and manage people-role assignments and transitions.
2. Edge Monitoring Agents:
$A^{*}_{e}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{e_{j}}\text{dist}(c_{j},A_{i}.\text{capabilities})$ (5)
These agents monitor dependencies between roles, where $c_{j}$ represents constraint types (temporal, spatial, etc.).
$\mathcal{MP}$ iteratively refines the network by validating workflow constraints and performance and replacing or adding agents, i.e., checking and adding nodes and edges.
4.4 Workflow Execution Framework
The workflow $\mathbf{W^{*}}$ requires a runtime system for execution. In this work, we evaluate $\mathbf{W^{*}}$ by entering $\mathbf{W^{*}}$ along with the problem statement into an LLM (DeepSeek and GPT4o). A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ must be performed manually.
5 Evaluating $\mathcal{MP}$’s Planning Capabilities
To assess $\mathcal{MP}$’s performance and adaptability, we adopted a dual-approach experimental structure. The first experiment uses the Traveling Salesperson Problem (TSP) to validate $\mathcal{MP}$’s optimization capabilities. The second involves the Thanksgiving Dinner Planning problem, showcasing $\mathcal{MP}$’s ability to handle complex, real-world challenges with cross-thread dependencies and dynamic adaptability.
Traveling Salesperson Problem (TSP)
The TSP experiment benchmarks $\mathcal{MP}$ against planners (Claude, DeepSeek, GPT4o) and their $\mathcal{MP}$ integrations. Metrics include: 1) solving quality, and 2) solution optimality. Detailed results are in Appendix E. The conclusion is that though the task is simple, $\mathcal{MP}$ helps improve results for monitoring execution, validating results, and suggesting algorithmic approaches.
Thanksgiving Dinner Planning
This problem, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraints and dependencies. It involves multiple agents performing interdependent tasks, representing a complex multi-agent scenario.
Planning performance is compared across three configurations: DeepSeek + $\mathcal{MP}$, GPT4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$. Metrics include:
[…]
where flexibility measures slack time incorporated to handle unexpected events.
5.1 Meta-Planning for Thanksgiving Event
Following Algorithm 1, $\mathcal{MP}$ generates workflows with:
  • Role nodes (e.g., cook, drivers, supervisor),
  • Explicit constraint edges (e.g., temporal, spatial, etc.),
  • Implicit constraint edges from common-sense analysis.
The planner monitors nodes and edges, enabling dynamic adjustments. Full specifications are in Appendix A.
Evaluation Scenarios
We test $\mathcal{MP}$ under:
  1. 1.
    Sequential Planning: Task executed as planned.
  2. 2.
    Reactive Planning: A 3-hour flight delay requiring task reallocations.
Meta-Planner Output
$\mathcal{MP}$ enhances planning by:
  • Identifying implicit constraints (e.g., luggage claim time, car rental delays),
  • Clarifying role dependencies,
  • Incorporating common-sense constraints (e.g., fatigue, social preferences),
In reactive planning, $\mathcal{MP}$ integrates an early information agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
5.2 Experimental Results
Sequential Planning Performance
With $\mathcal{MP}$’s enhanced workflow $\mathbf{W^{*}}$, all three LLMs successfully generated feasible solutions, a significant improvement over their previous failures with the original problem specification.
As illustrated in Table 6, DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations. While GPT4o eventually produced a valid solution in three iterations, it created suboptimal travel patterns by having Michael make separate trips. Claude’s solution, though feasible in two iterations, included unnecessary travel between pickup tasks. This experiment highlighted how $\mathcal{MP}$’s explicit constraint specification and common-sense augmentation enabled consistent performance improvement across different LLMs.
Table 6: Sequential Planning Performance. (# = iterations)
LLM # Notable Features
DeepSeek 2 Optimized airport wait time for James; balanced workload
GPT4o 3 Extra travel for Michael; suboptimal load balance
Claude 2 Unnecessary travel between pickup tasks
Reactive Planning Performance
GPT4o failed to produce a valid solution, violating three critical constraints without the ability to recognize these violations.
Reactive Planning Performance
The flight delay scenario revealed significant differences between LLMs’ capabilities. DeepSeek demonstrated superior spatial reasoning by routing Michael directly to the airport, an insight that should have come from $\mathcal{MP}$’s common-sense spatial reasoning. This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix A.9), which DeepSeek leveraged for immediate replanning.
In contrast, Claude produced two feasible plans but missed the 10:00 AM alert in $\mathbf{W^{R}}$, starting its schedule at 1:00 PM and missing opportunities for proactive actions like early Grandma pickup to free Sarah as a driver. GPT4o failed entirely, producing three constraint violations it couldn’t recognize, preventing further improvements.
Table 7: Reactive Planning Performance (Flight Delay)
LLM # Notable Features
DeepSeek 3 Smart routing of Michael directly to airport; efficient travel patterns
GPT4o Failed Failed to maintain critical constraints; unable to recover
Claude 3 Two valid plans with different trade-offs; longer wait times
6 Conclusions with Impact Statements
Current artificial intelligence systems excel at pattern matching but struggle with deliberate analytical reasoning required for complex planning. While Large Language Models (LLMs) demonstrate impressive capabilities, they face three critical limitations: inability to validate their own outputs (paralleling Gödel’s incompleteness theorems), inconsistent maintenance of constraints during planning, and inadequate integration of common-sense knowledge. Our experiments demonstrate that existing multi-agent systems, focused primarily on coordination, fail to address these limitations.
MACI (Multi-Agent Collaborative Intelligence) advances the state of the art through three key contributions:
  1. 1.
    Meta-Planning Architecture: MACI generates specialized planning workflows by:
    • Constructing role-node and dependency networks.
    • Identifying implicit constraints through common-sense analysis.
    • Establishing validation across distributed agents.
  2. 2.
    Workflow Enhancement: As demonstrated in our experiments:
    • Sequential Planning: Enhanced workflow $\mathbf{W^{*}}$ enabled all LLMs to generate feasible solutions.
    • Reactive Planning: Early information agent in $\mathbf{W^{R}}$ enabled proactive replanning.
    • Unexpected synergy: DeepSeek contributed spatial optimizations beyond MACI’s specifications.
  3. 3.
    Empirical Validation: Our Thanksgiving dinner planning case study shows:
    • Improving constraint satisfaction on LLMs.
    • Effective handling of crises through early detection.
    • Clear performance differences between LLMs in utilizing MACI’s capabilities.
MACI demonstrates that addressing LLM limitations requires more than simple agent coordination. It demands a structured meta-planning architecture that enables comprehensive constraint awareness and validation. The framework’s success in complex temporal planning, coupled with unexpected improvements from LLM contributions, suggests a promising direction for developing more capable AI planning systems.
Appendices
Appendix A Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
The problem statement remains consistent with Section 3, with $\mathbf{W^{*}}$ generated by $\mathcal{MP}$ to enhance constraints and dependencies. Planning performance is compared across four configurations: DeepSeek, GPT4o, DeepSeek + $\mathcal{MP}$, and GPT4o + $\mathcal{MP}$.
A.1 Phase 1: Network Construction
A.1.1 Node (Role) Specifications
First, meta-planner $\mathcal{MP}$ extracts roles ($\mathcal{N}$) with their required qualifications:
  • $n_{\text{cook}}$: capability to prepare dinner
  • $n_{\text{driver1}}$: capability to drive, pick up from airport
  • $n_{\text{driver2}}$: capability to drive, pick up grandma
  • $n_{\text{supervisor}}$: capability to monitor oven
A.1.2 Edge (Dependency) Specifications
Next, $\mathcal{MP}$ identifies dependencies ($\mathcal{E}$) between roles:
=$\mathcal{E}=\{e_{\text{temporal}},e_{\text{spatial}},e_{\text{safety}}\}$ (6)
The critical dependencies include:
  • $e_{\text{temporal}}$: - Turkey (4 hours) must finish by 6:00 PM - Side dishes (2 hours) must finish by 6:00 PM - Airport pickups must align with landing times
  • $e_{\text{spatial}}$: - Driver-passenger location matching - Travel time constraints between locations
  • $e_{\text{safety}}$: - Continuous oven supervision requirement
A.2 Phase 2: Agent Assignments
After constructing the network structure, $\mathcal{MP}$ selects and assigns agents to monitor both the roles and dependencies.
A.2.1 Node (Role) Agent Assignment
For each role, $\mathcal{MP}$ selects monitoring agents with the required capabilities:
$f_{\text{role}}:\mathcal{N}\rightarrow\mathbf{A}$ (7)
The role monitoring agents include:
  • Cook Monitor: Tracks cooking timeline, coordinates meal components
  • Driver Monitor: Validates driver availability
  • Supervisor Monitor: Ensures oven supervision
  • Resource Monitor: Manages vehicle assignments and actor schedules
A.2.2 Edge (Dependency) Agent Assignment
For the identified dependencies, $\mathcal{MP}$ assigns specialized monitoring agents:
$f_{\text{edge}}:\mathcal{E}\rightarrow\mathbf{A}$ (8)
Dependencies require these monitoring agents:
  • Temporal Agent: Manages timing constraints (cooking durations, travel times, arrival schedules)
  • Spatial Agent: Tracks location constraints (airport-home-grandma routes)
  • Safety Agent: Ensures oven supervision constraint remains satisfied
The resulting agent assignments create a complete monitoring system where:
  • Role agents track individual actor assignments and qualifications
  • Edge agents monitor interactions and dependencies between roles
  • All agents coordinate to maintain global constraint satisfaction
Table 8: Node (Role) Monitoring Agent Requirements
Agent Input Protocol Output Protocol
Cook
Monitor
Role: cook
Qualifications: skills
Time: prep and cook
Status: progress
Alerts: timing issues!
Updates: completed?
Driver
Monitor
Role: driver
Qs: license, rest
Where: current GPS
Status: availability
Alerts: fatigue warnings
Updates: new GPS
Supervisor
Monitor
Role: supervisor
Location: house
Duration: cover time
Status: covered?
Alerts: coverage gaps!
Updates: role transitions
Table 9: Edge (Dependency) Monitoring Agent Requirements
Agent Input Protocol Output Protocol
Temporal Start times
Durations
Deadlines
Buffer requirements
Schedule conflicts
Timing violations
Schedule updates
Spatial Locations
Routes
Travel times
Traffic conditions
Route violations
Location conflicts
Path updates
Safety Critical constraints
Resource states
Coverage requirements
Safety violations
Resource conflicts
Mitigation plans
A.2.3 Common Sense Constraint Analysis (Performed by an LLM)
A common sense agent identifies the following implicit constraints that can affect Thanksgiving dinner planning. This list is generated by Claude given the problem statement.
  • Physical Processing Times:
    • Airport luggage claim: 30 minutes
    • Car rental procedures: 30 minutes
    • Holiday traffic variations
    • Winter weather considerations
  • Human Factors:
    • Driver fatigue after long trips
    • Cooking preparation overhead
    • Multi-tasking limitations
    • Task switching delays
    • Required rest periods
  • Resource Dependencies:
    • Vehicle passenger capacity
    • Oven temperature management
    • Kitchen workspace limits
    • Shared resource coordination
  • Social Considerations:
    • Personal preferences for interactions
    • Family dynamics in assignments
    • Post-travel guest comfort
    • Host preparation requirements
Table 10: Complete Workflow Specification: Nodes, Edges, and Agent Assignments
Type Component Requirements Agent Protocol Dependencies
Node Components (Roles)
Node Cook Role
- Turkey (4hr)
- Side dishes (2hr)
- Kitchen management
- Time management
Input: schedule, resources, recipes
Output: task progress, completion
Monitor: kitchen_state() → status
Validate: cooking_constraints()
Connected to:
- Supervisor
- Resource edges
Node Driver1
(James/Michael)
- Valid license
- Airport navigation
- Car rental capable
- Rest state adequate
Input: flight times, routes
Output: location, ETA
Monitor: driver_state() → status
Validate: driver_constraints()
Connected to:
- Airport pickup
- Travel edges
Node Driver2
- Valid license
- Local navigation
- Availability window
- Rest state adequate
Input: pickup schedule, route
Output: location, ETA
Monitor: driver_state() → status
Validate: driver_constraints()
Connected to:
- Grandma pickup
- Travel edges
Node Supervisor
- Home presence
- Oven monitoring
- Safety awareness
- Time commitment
Input: cooking schedule, rules
Output: supervision status
Monitor: safety_state() → status
Validate: safety_constraints()
Connected to:
- Cook role
- Safety edges
Edge Components (Dependencies)
Edge Temporal - Schedule tracking
- Buffer management
- Sequence logic
- Critical path
Input: timestamps, durations
Output: schedule conflicts
Monitor: schedule_state() → alerts
Optimize: timeline_adjust()
Connects:
- All roles
- All activities
Edge Spatial - Location tracking
- Route optimization
- Traffic updates
- Distance constraints
Input: locations, routes
Output: travel updates
Monitor: location_state() → alerts
Optimize: route_adjust()
Connects:
- Drivers
- Locations
Edge Resource - Vehicle allocation
- Kitchen resources
- People availability
- Capacity limits
Input: resource demands
Output: allocation status
Monitor: resource_state() → alerts
Optimize: resource_adjust()
Connects:
- All roles
- All resources
Edge Safety - Oven monitoring
- Driving safety
- Food safety
- Critical rules
Input: safety requirements
Output: violation alerts
Monitor: safety_state() → alerts
Enforce: safety_rules()
Connects:
- All roles
- Critical tasks
A.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
The common sense constraints identified above require different verification approaches:
Agent-Required Information
These constraints need specialized agents to verify and quantify:
  • Airport Operations
    • United Airlines’ average luggage delivery time at BOS Terminal B
    • Terminal B to rental car center: shuttle schedule, walking options
    • Historical flight delay patterns for November at BOS
  • Weather and Traffic
    • Boston weather forecast for the event date
    • Historical traffic patterns on Thanksgiving days
    • Impact on airport-city-suburb travel times
  • Task Dependencies
    • Kitchen workflow analysis for parallel cooking tasks
    • Resource contention in meal preparation
    • Critical path identification in cooking timeline
Human Verification
Certain constraints require explicit human input to ensure that the planning process takes into account subtle interpersonal and individual considerations. These include:
  • Family Dynamics
    • Preferred pickup arrangements for Grandma (e.g., Grandma loves to have a grandson surprise her).
    • Optimal relationship-based task pairings.
    • Social comfort factors in assignments (e.g., Sarah and Grandma do not work together in the kitchen).
  • Personal Capabilities
    • Individual cooking experience levels.
    • Driver comfort with airport navigation.
    • Multi-tasking abilities of participants.
This separation ensures that agents focus on collecting quantifiable data while humans provide essential social and personal insights. $\mathcal{MP}$ can then integrate both types of information into the final workflow design.
A.3 Agent Requirements and Assignments
The $\mathcal{MP}$ requires two categories of agents. $\mathcal{MP}$ specifies their requirements in the protocol buffer format in Table 8 for the nodes and Table 9 for the edges, respectively.
Each agent must implement these protocols to participate in the workflow. The meta-planner selects agents from the pool based on their ability to satisfy these interface requirements. During execution, agents communicate through these standardized protocols while maintaining their specialized monitoring functions.
A.4 Monitoring Protocols and Dynamic Adjustments
The workflow monitoring operates through a hierarchical protocol system that enables both routine supervision and dynamic adjustments.
Basic Monitoring Protocol
Each agent maintains a continuous monitoring cycle:
[…]
For example, the temporal agent tracks schedule adherence:
$\Delta t=t_{\text{planned}}-t_{\text{actual}}\begin{cases}\text{normal}&\text{if }|\Delta t|<\text{buffer}\\ \text{warning}&\text{if }\text{buffer}\leq|\Delta t|<\tau\\ \text{violation}&\text{if }|\Delta t|\geq\text{ threshold }\tau\end{cases}$ (10)
Dynamic Adjustment Mechanism
When deviations occur, the system initiates a three-phase response:
1. Impact Assessment:
$\text{impact}(e)=\sum_{n\in\text{affected}(e)}\text{severity}(n)\times\text{urgency}(n)$ (11)
2. Solution Generation:
$S^{*}=\operatorname*{arg\,min}_{s\in\text{Solutions}}\{\text{cost}(s)|\text{feasible}(s)\}$ (12)
3. Coordination Protocol:
$\text{update}:(W_{\text{current}},S^{*})\rightarrow W_{\text{new}}$ (13)
For instance, if James’s flight is delayed:
  • Spatial agent detects arrival time change
  • Temporal agent calculates ripple effects
  • Role agents evaluate reassignment options
  • Safety agent verifies continued supervision coverage
The meta-planner $\mathcal{MP}$ coordinates these responses while maintaining global constraint satisfaction.
A.5 Integrated Workflow Network
Table 10 presents the resulting workflow network $\mathbf{W^{*}}$, which includes all nodes and edges, and their assigned agents and protocols.
  1. 1.
    Role Nodes:
    • Cook1: Sarah (primary) or Grandma (if at home) with 4-hour turkey + 2-hour sides
    • Driver1: James (after car rental) or Michael
    • Driver2: Available person after initial pickups
    • Supervisor: Must be present while turkey cooks
  2. 2.
    Dependencies:
    • Temporal: Verified airport processing + travel times
    • Spatial: Traveling routes with traffic consideration
    • Safety: Continuous oven supervision requirement
  3. 3.
    Agent Monitoring:
    • Temporal Agent: Schedules with verified buffer times
    • Spatial Agent: Real-time location and route mgmt.
    • Safety Agent: Role coverage for supervision
A.6 Agent Interaction Specifications
Please, see Table 11.
Table 11: Agent Interaction Protocols and State Transitions
Interaction Type Protocol State Transitions Validation Rules
Node-to-Node Interactions
[…]
Message: (task, duration, reqs.)
States: prep → cooking → comp.
Trigger: task_state_change()
Validate: coverage()
Alert: coverage_gap()
Driver1 $\leftrightarrow$ Driver2 Protocol: pickup_handoff()
Message: (location, time, passenger)
States: available → enroute → comp.
Trigger: location_change()
Validate: timing_feasible()
Alert: schedule_conflict()
Edge Agent Operations
Temporal Agent Protocol: schedule_monitor()
Message: (event, time, dependencies)
States: scheduled → active → comp.
Trigger: time_milestone()
Validate: timing_feasible()
Alert: delay_impact()
Spatial Agent Protocol: location_track()
Message: (actor, position, dest.)
States: idle → moving → arrived
Trigger: position_update()
Validate: route_feasible()
Alert: travel_delay()
A.7 New Problem Statement Revised with $\mathbf{W^{*}}$
Given the $\mathbf{W^{*}}$ generated by MACI’s meta-planner $\mathcal{MP}$, the Thanksgiving Dinner Planning problem statement stated at the beginning of this section is revised as follows:
Initial Setup:
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 1:00 PM Eastern time.
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James must rent a car after landing
  • Emily must be picked up from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven for safety
  • Side dishes require 2 hours of preparation, which can overlap with turkey
  • Travel time between home and Boston airport is one hour (one-way)
  • Travel between Boston airport and grandma home is one hour (one-way)
  • Travel between home and grandma home 30 minutes (one-way)
* New Dependencies:
  • The airport luggage pickup time after landing is 30 minutes.
  • Renting a car takes 30 minutes.
  • One person can simultaneously prepare turkey and side dishes.
  • Grandma prefers Michael to pick her up, provided that it does not cause the dinner time delay.
  • Grandma and Sarah prefer not to cook together in the kitchen.
  • Traffic congestion is not factored into current planning.
Planning Question Set:
  1. 1.
    All tasks and dependencies must be strictly observed in the plan, or the plan fails.
  2. 2.
    Dinner time is strictly at 6:00 PM, all tasks must be completed by then (redundancy).
  3. 3.
    Account for the idle time of each person.
  4. 4.
    The schedule consists of three columns: time, task, and assigned person(s).
A.8 Experiment #1: Sequential Planner
Once after the original plan was revised by $\mathcal{MP}$ to include more specific details, clarify ambiguous explicit constraints, and define implicit constraints, the performance of the three LLMs used in the experiment improved significantly. When the augmented plan $\mathbf{W^{*}}$ was input into DeepSeek, GPT4o, and Claude, each model successfully generated a feasible plan within two to three iterations. (The case study in Section 3 shows that DeepSeek was confusing and GPT4o repeatedly committed constraint violations.)
Table 12: DeepSeek’s Plan, Two Iterations
Time Task Assigned
1:00 PM James lands at Boston James
1:00–1:30 PM James picks up luggage James
1:30–2:00 PM James rents a car James
2:00 PM Turkey in oven (4 hours; requires monitoring) Sarah
2:00–3:00 PM James waits at airport (idle) James
2:30 PM Emily lands at Boston Emily
2:30–3:00 PM Emily waits for luggage Emily
3:00 PM James picks up Emily James
3:00 PM Michael arrives home Michael
3:00 PM Michael departs to Grandma Michael
3:30 PM Michael picks up Grandma Michael
3:30–4:00 PM Michael drives back home with Grandma Michael
3:00–4:00 PM James drives Emily home (airport to home: 1 hour) James
4:00 PM James and Emily home James
4:00 PM M. and Grandma home Michael
4:00–6:00 PM Sarah prepares side dishes Sarah
6:00 PM Thanksgiving dinner begins All
A.8.1 Results: DeepSeek Wins
Upon closer examination of the number of iterations required to produce a feasible plan, DeepSeek and Claude each required one revision (two iterations), while GPT4o required two revisions (three iterations). In terms of scheduling quality, measured by slack time, total driving distance, and load balance, DeepSeek (Table 12) outperformed both Claude (Table 14) and GPT4o (Table 13). DeepSeek optimized time and effort by scheduling James to wait at the airport for 30 minutes to pick up Emily. In contrast, Claude scheduled James to drive home and then return to the airport to pick up Emily, resulting in unnecessary travel. GPT4o assigned James to return home and scheduled Michael to first pick up Emily and then proceed to pick up Grandma, leading to a less balanced load. A better solution to reduce travel time would have been to schedule Michael to pick up Emily first and then drive with her to Grandma’s home to pick up Grandma, allowing all three to return home together. This adjustment would save 30 minutes of driving time and improve Grandma’s overall happiness to see both grandchildren.
Table 13: GPT4o’s Plan, Three Iterations
Time Task Assigned
1:00 PM Land at BOS Airport James
1:00-1:30 PM Luggage pickup James
1:30-2:00 PM Rent car James
2:00 PM Start turkey Sarah
2:00-3:00 PM Drive home James
2:30 PM Land at BOS Airport Emily
3:00 PM Arrive home Michael
3:00-4:00 PM Drive to airport, pick up Emily Michael
4:00-5:00 PM Return home with Emily Michael
5:00-5:30 PM Drive to Grandma’s Michael
5:30-6:00 PM Return with Grandma Michael
4:00-6:00 PM Prepare side dishes Sarah
6:00 PM Dinner served All
Table 14: Claude’s Plan, Two Iterations
Time Task Assigned
1:00 PM Land at BOS Airport James
1:00-1:30 PM Luggage pickup James
1:30-2:00 PM Rent car James
2:00 PM Start turkey Sarah
2:00-3:00 PM Drive home James
2:30 PM Land at BOS Airport Emily
3:00 PM Arrive home Michael
3:00-4:00 PM Drive to airport, pick up Emily James
4:00-5:00 PM Return home with Emily James
4:30-5:00 PM Drive to Grandma’s Michael
5:00-5:30 PM Return with Grandma Michael
4:00-6:00 PM Prepare side dishes Sarah
6:00 PM Dinner served All
A.8.2 Observations on Errors
Handling Long Dependencies
Complex scheduling problems reveal cascading errors when dependencies overlap. Key constraints, especially multi-factor dependencies, often get dropped during iterative problem-solving.
Reason: Cognitive load limits simultaneous constraint tracking, making exhaustive verification difficult in single passes.
Solution Framework:
  • Isolate and enumerate atomic task dependencies.
  • Verify global constraint satisfaction.
  • Implement systematic conflict resolution.
Stale Memory and Iterative Revisions
Iterative solutions can propagate errors due to partial constraint resets.
Reason: Over-reliance on previous solutions without full constraint re-evaluation leads to compounding errors.
Relation to Gödel’s Incompleteness:
  • Systems capable of arithmetic contain unprovable truths.
  • Similarly, inherited errors hinder consistent solutions.
  • Clean-state resets necessary for error prevention.
Implementation Strategy
Reset to baseline state for each iteration, fully re-evaluating all constraints.
Core Challenges:
  • Nested dependency management.
  • Residual error prevention.
  • Cross-iteration consistency.
A.9 Experiment #2: Reactive Planner for Flight Delay
At 10:00 AM Eastern time, Sarah is notified that James’s flight is delayed by three hours, with a new arrival time of 4:00 PM. Incorporating this unexpected delay, $\mathcal{MP}$ generates a reactive plan, $\mathbf{W^{R}}$.
Early Information Agent Addition
The meta-planner adds an early information agent to monitor upstream events:
$f_{\text{early}}:\mathcal{E}_{\text{upstream}}\rightarrow\text{alerts}$ (14)
The agent’s protocol is defined as:
Table 15: Early Information Agent Specification
Component Flight Monitor Impact Analyzer
Input Flight status, departure logs, weather Alert details, workflow dependencies
Output Alert(event, severity, delay) Replan(affected_nodes, time_window)
This addition allows the workflow to initiate replanning at the earliest possible moment when upstream changes occur, significantly enhancing the system’s proactive planning capability. Since none of the planned elements have been executed, this reactive planning effectively functions as proactive planning.
In this experiment, the problem statement remains unchanged apart from James’s updated arrival time.
Initial Setup (Updated at 10:00 AM):
  • Mom (Sarah) is hosting Thanksgiving dinner at 6:00 PM in Boston. The following family members are traveling:
  • Dad (James) flying from San Francisco, landing at 4:00 PM Eastern time [UPDATED].
  • Sister (Emily) flying from Chicago, landing at 2:30 PM
  • Brother (Michael) driving from New York, estimated arrival 3:00 PM at home
  • Grandma is healthy and needs to be picked up from her home in suburban Boston
Critical Dependencies:
  • James must rent a car after landing
  • Emily must be picked up from airport, no other transportation options are allowed
  • Turkey needs 4 hours to cook, someone must be in the house once turkey is in oven for safety
  • Side dishes require 2 hours of preparation, which can overlap with turkey
  • Travel time between home and Boston airport is one hour (one-way)
  • Travel between Boston airport and grandma home is one hour (one-way)
  • Travel between home and grandma home 30 minutes (one-way)
* New Dependencies:
  • The airport luggage pickup time after landing is 30 minutes.
  • Renting a car takes 30 minutes.
  • One person can simultaneously prepare turkey and side dishes.
  • Grandma prefers Michael to pick her up, provided that it does not cause the dinner time delay.
  • Grandma and Sarah prefer not to cook together in the kitchen.
  • Traffic congestion is not factored into current planning.
Planning Question Set:
  1. 1.
    All tasks and dependencies must be strictly observed in the plan, or the plan fails.
  2. 2.
    Dinner time is strictly at 6:00 PM, all tasks must be completed by then (redundancy).
  3. 3.
    Account for the idle time of each person.
  4. 4.
    The schedule consists of three columns: time, task, and assigned person(s).
A.9.1 Results: DeepSeek Wins
None of the LLMs cannot react appropriately to this new event without clearing their context buffers. As explained in Appendix A.8.2, this limitation is evident. The key takeaway is that for future runtime frameworks, we must ensure infrastructure support for selectively invalidating stale constraints. If a workflow is already in execution, completed steps and assignments cannot be erased or altered. For example, in a stock-market investment plan, when pertinent news arrives, $\mathcal{MP}$ cannot revert completed nodes or resolved dependencies in $\mathbf{W^{R}}$. For now, we treat the reactive plan as a new plan, given that no steps have been realized in the real world by 10:00 AM.
Table 16 presents GPT4o’s plan. There are three severe constraint violations. Unfortunately, when asked to identify violations, it answers none. Therefore, $\mathcal{MP}$ is stuck without a feasible plan.
Table 16: GPT4o’s Infeasible Plan. Fail to proceed.
Time Task Assigned
10 - 2:00 PM Prep side dishes (2 hours, overlaps with turkey cooking later) Sarah
X2:30 - 3:00 PM Pick up Emily from the airport Sarah
X3:00 - 4:00 PM Prep turkey and place it in the oven (4-hour cooking time) Sarah
3:00 - 3:30 PM Michael drives to pick up Grandma Michael
3:30 - 4:00 PM Drive Grandma home Michael
4:00 - 4:30 PM James lands and gets luggage James
4:30 - 5:00 PM James rents a car James
X5:00 - 5:30 PM James drives home James
5:00 - 6:00 PM Set the table and clean kitchen Emily
5:30 - 6:00 PM Michael helps Grandma settle in and assists with final prep Michael
Table 17 depicts Claude’s plan. It violated a couple of constraints in the first two attempts, but these were minor. For instance, in the second trial, it planned for Michael’s round trip to Grandma’s home to take 30 minutes. However, the key is that Claude can recognize its own error and make corrections in the next iteration.
When asked to produce an alternate plan to reduce wait time and improve load balancing, as Michael can suffer from severe fatigue, an implicit constraint, Claude generates another feasible plan in Table 18. In this plan, James picks Emily instead of Michael. Emily has to wait for James’ availability for two hours at the airport.
There are clearly other alternatives to improve the schedule and eliminate Emily’s wait time, but none of the LLMs can figure that out. For example, using the time between 10:00 am and 1:00 pm, Sarah could have picked up Grandma and assigned her to cook, allowing Sarah to be available as a driver.
Table 17: Claude’s Reactive Plan #1, Three Iterations
Time Task Assigned
13:00 Start turkey in oven Sarah
14:00 Start side dishes Sarah
14:30 Land at airport Emily
15:00 Arrive home from NY Michael
15:00 Leave for airport Michael
16:00 Land at airport James
16:00 Pick up Emily Michael
17:00 Arrive home with Emily Michael
17:00 Leave for Grandma Michael
18:00 Arrive home with Grandma Michael
18:00 Arrive home James
18:00 Dinner starts All
Table 18: Claude’s Reactive Plan #2. Michael can rest.
Time Task Person
13:00 Start turkey Sarah
14:00 Start side dishes Sarah
14:30 Land at airport Emily
15:00 Arrive from NY Michael
16:00 Land at airport James
16:30 Leave for Grandma Michael
16:30 Get rental car James
17:00 Pick up Emily James
17:00 Pick up Grandma Michael
17:30 Return home with Grandma Michael
18:00 Arrive with Emily James
18:00 Dinner starts All
DeepSeek offers a clever alternative by routing Michael directly to the Boston airport without stopping at home first. This is a pleasant common-sense inference that the other two LLMs failed to include themselves. (This was supposed to be provided by $\mathcal{MP}$’s common-sense spatial reasoning, but it did not.)
However, Michael could drive to Grandma’s home after picking up Emily. This schedule not only saves 30 minutes but also makes Grandma happy by allowing her to surprisingly see two grandchildren at the same time.
Table 19: DeepSeek’s Reactive Plan. Three Iterations. Routining Michael directly to BOS is smart.
Time Task Assigned
10:00 AM Michael departs New York for Boston Airport (4-hour drive). Michael
2:00 PM Start cooking turkey Sarah
2:30 PM Emily lands at Boston Emily
3:00 PM Emily gets her luggage Emily
3:00 PM Michael arrives at Logan airport, picks up Emily. Michael
3:00–4:00 PM Drive Emily home Michael
4:00 PM Michael departs for Grandma Michael
4:00 PM James lands at Boston Airport James
4:00–4:30 PM James picks up luggage James
4:30–5:00 PM James rents car (30 minutes). James
4:30 PM Michael arrives at Grandma’s Michael
5:00 PM Michael & Grandma home. Grandma
5:00–6:00 PM James drives home from BOS James
4:00–6:00 PM Sarah prepares side dishes (overlaps with turkey). Sarah
6:00 PM James arrives home. Dinner served. All
A.9.2 Observations on Errors
The initial meta-planner failed to recognize a critical opportunity in early information detection. A flight delay from SFO to BOS becomes known at departure time (10:00 AM EST) rather than arrival time (1:00 PM EST). An early information agent could enable replanning three hours sooner by monitoring flight departures. To remedy this oversight, the meta-planner adds an early information agent specification, detailed in Table 15. DeepSeek was aware of this alert in a timely manner, but Claude was not.
A.10 Conclusion
Our concluding remark is that we may not be able to rely on LLMs alone to cover all constraints and react promptly to various alerts. This reinforces that the MACI architecture is on the right path to address all the aforementioned limitations of LLMs, some of which cannot be rectified.
Appendix B Validation and Recovery Protocols
The validation protocol implements a multi-stage process for ensuring state consistency. When any agent proposes a state change, the validation agent initiates a sequence of checks:
$\text{validate}(s_{t}\rightarrow s_{t+1})=\begin{cases}\text{true}&\text{if all checks pass}\\ \text{false}&\text{if any check fails}\end{cases}\vspace{-.1in}$ (15)
The protocol begins with pre-validation. Before a state transition starts, the validation agent queries relevant agents about preconditions. For a travel booking, temporal agent verifies the proposed times fit within existing schedules. Spatial agent confirms the physical feasibility of movements between locations. Role agent checks if all actors can perform their assigned functions.
During the transition, the protocol maintains atomic operations. The validation agent tracks changes across all state dimensions, ensuring partial updates cannot create inconsistent states. If the temporal agent approves a flight time but the resource agent finds insufficient seats, the entire transition fails and rolls back.
Post-validation examines the resulting state. The validation agent verifies that all constraints remain satisfied after the change. Common sense agent reviews the new state for practical issues that formal checks might miss. Strategy agent confirms the transition aligns with overall planning objectives.
When validation fails, the protocol triggers a structured recovery process:
$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (16)
Recovery begins by logging the failure cause and violated constraints. The strategy agent then works with domain agents to generate alternative proposals that satisfy the constraints. This might involve relaxing non-critical constraints or exploring different approaches to meet the planning objectives.
B.1 Operations Research Techniques in Validation Protocols
The validation protocols described above align closely with established methods in operations research (OR). Some relevant techniques include:
  • Constraint Programming (CP): Focuses on solving combinatorial problems by enforcing constraints, ensuring consistency across dimensions such as temporal, spatial, and resource availability (Rossi et al., 2006).
  • Mixed-Integer Linear Programming (MILP): Optimizes decision variables subject to linear constraints and objective functions, often used in scheduling and resource allocation (Wolsey & Nemhauser, 1998).
  • Network Flow Algorithms: Validates feasibility and optimizes flows in networks by ensuring capacity, timing, and availability constraints are satisfied (Ahuja et al., 1993).
  • Dynamic Programming (DP): Decomposes problems into sequential subproblems, useful for validating processes like inter-terminal walking or luggage claiming time (Bellman, 1957).
  • Monte Carlo Simulation: Simulates scenarios to validate feasibility and robustness under uncertainty (Metropolis & Ulam, 1949).
  • Robust Optimization: Focuses on solutions that remain feasible under uncertainty, ensuring plans adapt to disruptions (Ben-Tal & Nemirovski, 2009).
Integration in Multi-Agent Systems
The validation agent also employs techniques from multi-agent systems, such as:
  • Blackboard Systems: A shared workspace for collaborative validation by different agents, ensuring global consistency (England & Engelmore, 1987).
  • Consensus Protocols: Used for distributed validation, where agents negotiate to ensure all constraints are met (Ren & Beard, 2005).
By combining these OR techniques with agent-based systems, the validation protocol ensures comprehensive and adaptive checks for workflow consistency. Future work can explore integrating heuristic methods, such as genetic algorithms or simulated annealing, to further enhance recovery processes.
Appendix C $\mathsf{MACI}$ Additional Design Considerations
C.1 Cross-Domain Generalization
While the state space dimensions—Who, Where, When, What, and Why—are broad enough to cover diverse domains, additional customization may be required for unique applications. This section examines how MACI generalizes across domains like financial planning, healthcare logistics, and supply chain optimization. The travel planning example is illustrative, emphasizing how MACI dynamically adapts state spaces and agents to domain-specific requirements.
C.2 Dynamic Agent Registration and Evolution
This section explores how agents are dynamically developed, trained, and validated for new tasks. It discusses mechanisms for evaluating new agents and integrating them into the repository without retraining the entire system, ensuring scalability and adaptability.
C.3 Scalability and Resource Efficiency
As the number of agents and task complexity grows, MACI employs strategies to manage communication overhead and optimize agent interactions. This section details techniques for clustering agents and hierarchical coordination to maintain scalability.
C.4 Empirical Evaluation Across Domains
To demonstrate MACI’s adaptability, this section presents empirical results from applying the framework to multiple domains. Examples include financial portfolio management, urban traffic planning, and hospital resource allocation, highlighting MACI’s advantages over state-of-the-art systems.
C.5 Challenges and Future Directions
While MACI addresses many limitations of LLM-based planning, challenges remain in real-time coordination, implicit knowledge integration, and robust recovery mechanisms. This section proposes future research directions to enhance MACI’s performance and applicability to novel tasks.
Appendix D Additional Tables and Figures
D.1 Notation
Table 20 presents all symbols used throughout this paper.
Table 20: Symbol Definitions
Symbol Definition Symbol Definition
Basic Sets
$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
$\mathcal{C_{E}}$ Explicit constraints $\mathcal{C_{I}}$ Implicit constraints
$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
Workflow Components
$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
$n_{i}$ Individual role $e_{ij}$ Role dependency
Agent Functions
$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
$V(\cdot)$ Validation function $\text{dist}(\cdot)$ Capability distance
$\text{map}_{\text{role}}$ Role extraction $\text{map}_{\text{edge}}$ Dependency extraction
Optimization
$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
D.2 Example Common Agents
  1. 1.
    Role Manager Agent (Who): Tracks actors, their roles, and their associated constraints, ensuring that all role-based requirements are satisfied.
  2. 2.
    Spatial Agent (Where): Manages geographic and location-based constraints, verifying transitions between physical or virtual locations.
  3. 3.
    Temporal Agent (When): Handles scheduling, timing, and deadlines, ensuring alignment with temporal constraints.
  4. 4.
    Resource Agent (What): Tracks real-world resources such as tools, vehicles, or financial instruments, managing capacity, availability, and associated costs.
  5. 5.
    Reasoning and Explanation Agent (Why): Maintains the rationale behind decisions, dependencies, and alternative plans, enabling consistent alignment with objectives and providing explanations for outcomes.
  6. 6.
    Common Sense Agent: Identifies implicit constraints, integrates practical knowledge, and ensures plans align with real-world considerations.
  7. 7.
    Constraint Validation Agent: Ensures that all constraints are satisfied and that proposed plans remain feasible.
  8. 8.
    Plan Evaluation Agent: Assesses the effectiveness of plans against predefined metrics and objectives.
  9. 9.
    What-If Testing Agent: Evaluates plan robustness by simulating alternative scenarios and analyzing their impact.
  10. 10.
    Compliance and Safety Agent: Monitors adherence to safety standards, ethical principles, and regulatory frameworks.
Appendix E Traveling Salesman Problem Experiment
E.1 General Problem Specification
The TSP requires finding the shortest possible route visiting N locations exactly once, returning to the start:
  • Inputs: N locations, distance matrix D[N][N]
  • Output: Optimal tour T minimizing total distance
  • Constraints: Each location visited once, return to start
Computational Complexity - Brute Force
For N locations:
  • Number of possible tours = (N-1)!/2
  • Time complexity = O(N!)
  • Space complexity = O(N2)
Solution Methods
  1. 1.
    Exact Methods: Representative methods are Branch and Bound (Land & Doig, 1960), Dynamic Programming (Bellman, 1962), and Integer Linear Programming (Dantzig et al., 1954).
  2. 2.
    Heuristics: Methods include Nearest Neighbor, Insertion Methods, and Christofides Algorithm (3/2-approximation) (Christofides, 1976).
  3. 3.
    Meta-heuristics: This category includes Genetic Algorithms (Holland, 1992), Simulated Annealing (Kirkpatrick et al., 1983), and Ant Colony Optimization (Dorigo & Stützle, 2004).
E.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
Node Components (N)
For TSP with n locations:
$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (17)
    $n_{\text{route}}$: Route generation role
    $n_{\text{dist}}$: Distance calculation role
    $n_{\text{valid}}$: Solution validation role
Edge Dependencies (E)
$E=\{e_{\text{spatial}},e_{\text{sequence}},e_{\text{complete}}\}\text{ where}$ (18)
    $e_{\text{spatial}}$: Distance constraints between locations
    $e_{\text{sequence}}$: Visit order constraints
    $e_{\text{complete}}$: Tour completion requirements
Agent Assignments
Node Agents ($A_{n}$):
  • Route Generation Agent: Generates candidate tours
  • Distance Calculator Agent: Computes tour lengths
  • Solution Validator Agent: Verifies tour validity
Edge Agents ($A_{e}$):
  • Spatial Constraint Agent: Monitors distance feasibility
  • Sequence Monitor Agent: Ensures valid visit order
  • Completion Checker Agent: Verifies tour completeness
Algorithm Selection
Based on the size of the problem, an algorithm is selected to balance performance trade-offs and mitigate the exponential computational cost of the brute-force method.
Validation Function
$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$ (19)
E.3 Experiments, From Small to Large N
  1. 1.
    N=5: Establish ground truth via brute force
  2. 2.
    N=10,20,100: Test LLM performance degradation
  3. 3.
    Metrics:
    • Solution quality vs optimal
    • Computation attempts before giving up
    • Error recognition capability
E.3.1 Small Campus Tour (N=5)
Plan an optimal route for campus tour guide visiting 5 key locations:
  • A: Admissions Office (start/end)
  • B: Library
  • C: Student Center
  • D: Science Building
  • E: Sports Complex
Distance Matrix (minutes)
$D=\begin{bmatrix}0&5&8&4&7\\ 5&0&6&3&8\\ 8&6&0&5&4\\ 4&3&5&0&6\\ 7&8&4&6&0\end{bmatrix}$ (20)
Constraints
  • Tour starts/ends at Admissions (A)
  • Each location visited exactly once
  • Total possible routes: (5-1)!/2 = 12
  • Optimal solution can be verified by hand
3.1.1. $\mathbf{W^{*}}$ Workflow Components
Node Components (N)
  • $n_{\text{route}}$: Generates permutations starting/ending at A.
  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
Edge Dependencies (E)
  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
  • $e_{\text{sequence}}$: Ensures visit order consistency.
  • $e_{\text{complete}}$: Validates all 5 locations are visited.
Agent Assignments
  • Node Agents:
    • Route Generation Agent (handles $n_{\text{route}}$)
    • Distance Calculator Agent (handles $n_{\text{dist}}$)
    • Solution Validator Agent (handles $n_{\text{valid}}$)
  • Edge Agents:
    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
    • Completion Checker Agent (enforces $e_{\text{complete}}$)
Selected Algorithm
Brute-force.
Validation Function
$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$
3.1.2. Solution Steps
Step 1: Problem Parsing
  • Input: 5 locations with distance matrix $D$.
  • Initialize node/edge agents and constraints.
Step 2: Route Generation ($n_{\text{route}}$ Agent)
  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
  • Example permutations:
    • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
    • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
  • Compute total time for each route using $D$.
Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
  • Check all routes for:
    • Start/end at A (e.g., invalid route: $A\rightarrow B\rightarrow C\rightarrow D\rightarrow E\rightarrow B$).
    • No duplicate visits.
Step 5: Edge Agent Validation
  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
  • Sequence Monitor Agent: Confirm no backtracking (e.g., $B\rightarrow D$ allowed; $D\rightarrow B$ invalid unless part of loop).
Step 6: Apply Validation Function
  • Assign $V=-\infty$ to invalid routes.
[…]
  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
3.1.3 Solution
Optimal tour time: 242424 minutes, achieved by three routes:
  • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
  • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
  • $A\rightarrow E\rightarrow C\rightarrow B\rightarrow D\rightarrow A$
E.3.2 Large Campus Tour (N=10)
Plan an optimal route for a guided tour through 10 locations:
  • Locations: A (Admissions), B (Library), C (Student Center), …, J (Sports Complex)
  • Distance Matrix: Asymmetric travel times (minutes)
$D=\begin{pmatrix}0&12&8&15&9&14&7&11&10&6\\ 10&0&7&14&6&16&9&13&5&8\\ 9&5&0&11&8&12&10&7&15&4\\ 14&8&12&0&10&9&13&6&11&7\\ 7&13&6&9&0&8&5&12&14&10\\ 11&9&15&8&12&0&7&10&13&5\\ 5&7&10&6&11&9&0&8&12&15\\ 8&14&4&10&7&13&6&0&9&11\\ 12&6&9&7&15&10&8&5&0&14\\ 9&10&7&13&5&11&14&8&12&0\end{pmatrix}$
Table 21: Comparison of Planners and Their Performance Characteristics. $\mathcal{MP}$ provides validation to improve accuracy.
Planner Best Results Algorithm Iters. Advantages Limitations
Claude $92\rightarrow 66$ mins Nearest Neighbor 3 Efficiently implements greedy heuristic approach Makes data reading errors, compromising solution accuracy
GPT4o $75\rightarrow 68$ mins Genetic 3 Identifies effective termination conditions Unable to implement exact algorithms like Held-Karp
DeepSeek $60$ mins Held-Karp 1 Implements optimal algorithm correctly None observed for this problem size
$\mathcal{MP}$ + Claude $66\rightarrow 60$ mins Ant Colony Optimization 2 Provides validation and suggests iteration increases for improvement Requires external guidance for algorithm selection and parameter tuning
MP + GPT4o $60$ mins Ant Colony Optimization 1 Achieves optimal solution with precise execution Requires more computational resources with larger ant population and iteration count
MP + DeepSeek $60$ mins Ant Colony Optimization 1 Combines efficient algorithm selection with optimal parameter tuning None significant for given problem
Algorithm Selection
Based on the size of the problem, $\mathcal{MP}$ selected the Ant Colony Optimization (Dorigo & Stützle, 2004) algorithm to achieve at least a 4x speedup. For $N\geq 10$, an approximate method is recommended.
3.2.1. ACO Method
Parameters
  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
The termination criteria can be modified to stop the algorithm if no meaningful improvement is observed after $k$ consecutive iterations.
Algorithm
1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
2:  for 50 iterations do
3:     for all 100 ants do
4:        Build tour using $P_{ij}=\frac{[\tau_{ij}]^{1}[\eta_{ij}]^{2}}{\sum[\tau_{ik}]^{1}[\eta_{ik}]^{2}}$
5:        Record tour length $L_{k}$
6:     end for
7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
9:     Track best tour
10:  end for
3.2.2. Performance Comparison
The goalpost is the optimal time of 606060 minutes. Table 21 compares six different configurations, and three of the six achieve the optimal answer. Although TSP is a relatively simple scheduling problem with just one actor and no parallel execution, the benefit of having $\mathcal{MP}$ to validate results is still helpful to Glaude and GPT4o.
When asked to solve the problem without $\mathcal{MP}$, Glaude and GPT4o initially chose brute force, then switched to an approximation method without thorough deliberation (or perhaps they did, but did not output their reasoning process). However, DeekSeek picked Held-Karp, a computationally expensive method, even more expensive than brute force, arguing that the absolute computation time for $N=10$ is only $0.2$ seconds. $\mathcal{MP}$ was more deliberate, opting for brute force when $N=5$ and ACO for $N=10$.
E.4 TSP Experiment Conclusion
This simple task demonstrates that $\mathcal{MP}$ can be valuable for monitoring the execution process, validating the correctness of intermediate results, and suggesting more efficient algorithmic approaches.

### v2

@@
-MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning
+MACI: Multi-Agent Collaborative Intelligence for Adaptive Reasoning and Temporal Planning
 Abstract
-Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities beyond the pattern-matching strengths of LLMs.
-LLMs struggle with planning tasks because of their reliance on associative reasoning, inability to self-verify, and inconsistent constraint awareness.
-We propose Multi-Agent Collaborative Intelligence (MACI), a framework centered on a meta-planner (MP) that orchestrates multiple agents to generate planner templates that define roles and constraints.
-These planners produce actionable workflows of role nodes and dependency constraints, enabling advanced temporal reasoning and adaptability.
-MACI’s three-tier architecture includes a meta-planning module for planner construction, common agents for general reasoning, and specialized agents for domain expertise.
-By decoupling planning from validation, it overcomes key LLM limitations.
-Evaluations demonstrate MACI’s effective constraint satisfaction, conflict detection, and reasoning, positioning it as a robust solution for complex reasoning and planning tasks.
+Artificial intelligence requires deliberate reasoning, temporal awareness, and effective constraint management—capabilities traditional LLMs often lack due to their reliance on pattern matching, limited self-verification, and inconsistent constraint handling.
+We introduce Multi-Agent Collaborative Intelligence (MACI), a framework comprising three key components: 1) a meta-planner (MP) that identifies, formulates, and refines all roles and constraints of a task (e.g., wedding planning) while generating a dependency graph, with common-sense augmentation to ensure realistic and practical constraints; 2) a collection of agents to facilitate planning and address task-specific requirements; and 3) a run-time monitor that manages plan adjustments as needed.
+By decoupling planning from validation, maintaining minimal agent context, and integrating common-sense reasoning, MACI overcomes the aforementioned limitations and demonstrates robust performance in two scheduling problems.
 Machine Learning, ICML
 1 Introduction
-The development of artificial general intelligence requires a shift from pattern matching to deliberate analytical reasoning.
-Large Language Models (LLMs) excel at pattern matching but lack deliberative reasoning required for planning tasks (Kahneman, 2011).
-These capabilities become crucial for tasks requiring sustained attention, awareness of constraints, and reasoning about future states.
+Advancing artificial intelligence requires capabilities beyond pattern matching.
+To tackle complex real-world tasks, AI must exhibit deliberate reasoning, temporal awareness, and effective constraint management.
+While Large Language Models (LLMs) excel at pattern recognition, they face significant challenges in planning tasks that demand sustained attention, comprehensive constraint awareness, and reasoning across both past and future temporal states (Kahneman, 2011).
 1.1 Limitations of LLMs in Planning
-Our evaluation reveals three critical limitations in current LLMs that prevent effective planning.
-Metacognitive Limitations in Self-verification
-LLM limitations in self-verification extend beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967).
-Though formal systems cannot prove their own consistency within their framework, LLMs face greater challenges due to their probabilistic nature and lack of logical foundations (detailed arguments in Section 2).
-This limitation provides theoretical justification for distributed validation, where specialized agents perform independent verification of planning processes.
-Attention-based Processing Bias
-In complex planning scenarios, LLMs demonstrate a form of cognitive tunneling where recent context dominates earlier-established constraints.
-For example, when planning a multi-stop journey, an LLM might focus on optimizing the final leg’s arrival time while losing sight of earlier constraints about vehicle availability or required rest periods.
-This attention bias leads to locally optimized but globally invalid solutions, undermining the overall feasibility of the plan.
-The inability to maintain consistent awareness of all constraints throughout the planning process represents a fundamental limitation in current architectures.
-Common Sense Knowledge Integration
-LLMs often overlook practical constraints that humans naturally consider.
-In travel planning, they might ignore factors such as airport security wait time or luggage claim delays.
-These omissions are the result of an inadequate integration of common sense knowledge into the planning process.
-Although humans automatically incorporate such practical considerations, LLMs require explicit specification of these constraints, often leading to incomplete or impractical plans.
-Table 1: Comparison of Validation Limitations and Solutions
-Gödel’s Incomp. Theorems LLM Validation Limitations MACI’s Solution
-A formal system cannot prove its own consistency LLMs cannot reliably validate their own outputs due to probabilistic foundations MACI uses external validation agents to bypass self-referential loops
-Consistency proofs require a “higher” system Validation requires external frameworks (e.g., logic engines, human oversight) Distributed validation acts as the “higher" system for LLM outputs
-Self-reference leads to paradox or incompleteness LLMs hallucinate or contradict themselves when critiquing their own outputs Meta-planning separates planning and validation into distinct agent roles
+LLMs reveal three limitations that fundamentally undermine their effectiveness in complex planning scenarios:
+1.
+Lack of Self-Verification.
+LLMs struggle with validating their own output, a problem that extends beyond Gödel’s incompleteness theorems for formal systems (Gödel, 1967).
+Their probabilistic nature and lack of logical foundations create significant barriers to self-assessment.
+This intrinsic limitation means LLMs cannot reliably detect errors or inconsistencies in their generated content, necessitating external mechanisms to validate and refine their output.
+2.
+Attention Bias and Constraint Drift.
+In complex scenarios, LLMs demonstrate a critical cognitive limitation known as cognitive tunneling.
+This phenomenon occurs when recently provided context dominates and progressively erodes earlier-established constraints.
+When planning a multi-leg journey, for instance, an LLM might optimize the final travel segment while completely neglecting crucial earlier constraints such as vehicle availability or required rest periods.
+This bias toward local optimization fundamentally undermines the global feasibility of generated plans.
+3.
+Lack of Common Sense Integration.
+LLMs often overlook practical constraints that humans intuitively consider.
+This deficiency becomes particularly evident in domains that require real-world experience and understanding.
+In travel planning, an LLM might generate a route without accounting for airport security processing times.
+In logistics, it may create schedules that ignore resource availability and preparation windows.
+Without explicit, granular specifications, these models produce plans that appear superficially coherent but remain impractical.
 1.2 The MACI Framework
-To overcome the inherent limitations of single-model approaches in planning tasks, we introduce the Multi-Agent Collaborative Intelligence (MACI) framework.
-At the core of MACI lies the meta-planner (MP), a high-level orchestration mechanism that collaborates with multiple agents to construct a specialized planner for temporal reasoning and scheduling.
-By enabling agents to work collectively at the meta-level, MACI ensures robust and adaptive planning across complex domains.
-Meta-Planning Module
-The meta-planning module serves as the cornerstone of MACI, analyzing task requirements and dynamically generating a planner specifically tailored to address complex challenges.
-This planner, in turn, produces a workflow composed of nodes representing roles (e.g., cook, driver, supervisor) and dependencies representing constraints (e.g., temporal, spatial, resource).
-Unlike traditional multi-agent systems, MACI’s meta-planner orchestrates a dynamic network of agents, validates their interactions, and enforces protocols to ensure comprehensive planning and execution.
-This layered approach—planning to generate a planner that produces actionable workflows—positions MACI as a forward-looking solution to real-world planning problems.
-Common Agents
-Common agents form the foundation of MACI’s general-purpose planning capabilities.
-They address critical tasks such as practical reasoning, constraint validation, and robustness analysis.
-For example, the Common Sense Integration agent identifies implicit constraints, the Constraint Validation agent ensures solution feasibility, and the Plan Evaluation agent measures performance metrics.
-These agents collectively maintain both explicit and implicit constraints while ensuring compliance and safety across domains.
-Specialized Agents
-Specialized agents extend MACI’s capabilities by addressing domain-specific requirements.
-Instead of overburdening a single planner, MACI deploys focused agents for tasks such as travel scheduling, financial risk assessment, or logistics optimization.
-These agents integrate domain expertise with the foundational work of common agents, ensuring precise and effective solutions for specialized planning challenges.
-1.3 Contributions
-This work introduces MACI as a novel approach to collaborative planning.
-Preliminary evaluations in travel planning scenarios demonstrate MACI’s superiority in constraint satisfaction, conflict detection, and practical reasoning compared to single-LLM approaches.
-Key contributions of this paper include:
-  1. 1.
-    Identifying limitations of single-LLM approaches through systematic experimentation.
-  2. 2.
-    Developing a three-tier architecture for distributed planning and validation.
-  3. 3.
-    Designing MACI with common and specialized agents for constraint management and reasoning.
-  4. 4.
-    Demonstrating MACI’s effectiveness in travel planning scenarios, achieving significant improvements over single-LLM systems.
-The remainder of this paper is organized into five sections.
-Section 2 reviews related MAS frameworks and explains why they fall short for the task at hand.
-Section 3 examines the limitations of LLMs through case studies.
-Section 4 details the specifications of the MACI framework.
-Section 5 evaluates MACI’s performance in comparison to state-of-the-art LLMs.
-Finally, Section 6 discusses implications for future research and alternative perspectives.
+To address these limitations, we propose Multi-Agent Collaborative Intelligence (MACI), a framework designed to enhance reasoning and planning through a multi-component architecture.
+MACI introduces three core components:
+1.
+Meta-Planner (MP).
+The meta-planner serves as the central orchestration mechanism in MACI.
+It analyzes task requirements, identifies roles and constraints, and dynamically generates a dependency graph (or workflow template) tailored to the task.
+This template includes actionable workflows with nodes representing roles (e.g., cook, driver, supervisor) and edges representing dependencies (e.g., temporal, spatial, or resource constraints).
+The incorporation of common sense augmentation into the metaplanner ensures that the generated plans are realistic, comprehensive, and aligned with practical constraints.
+2. Common and Task-specific Agents. MACI employs two types of agents to execute the generated plans:
+    Common Agents: These agents handle general-purpose tasks, including constraint validation, practical reasoning, and performance evaluation.
+For instance, a Common Sense Integration agent identifies implicit constraints, while a Constraint Validation agent ensures feasibility and compliance with the task’s requirements.
+    Task-specific Agents: These agents cater to domain-specific requirements, including task-dependent data and knowledge augmentation, selection of the most effective planning algorithms, safety and ethics assessment, and emergency response optimization.
+By integrating domain expertise, they extend the capabilities of common agents, enabling MACI to address specialized planning challenges with precision and adaptability.
+3.
+Run-Time Monitor.
+The run-time monitor handles real-time adjustments to the static plan in response to unexpected changes, such as resource delays, environmental disruptions, or evolving task requirements.
+This component ensures adaptability and robustness by:
+    Monitoring plan execution to detect deviations.
+    Activating emergency agents to revise dependencies, reassign roles, or dynamically adjust constraints.
+    Communicating updates to affected agents to maintain coherence throughout the workflow.
+1.3 Summary: How MACI Addresses LLM Limitations
+With its multi-component architecture, MACI directly addresses the three critical limitations of LLMs in planning:
+1.
+Lack of Self-Verification.
+MACI separates planning from validation, employing independent agents for validation.
+These agents operate without shared memory or interference, ensuring external verification of outputs and mitigating the risks of self-referential errors.
+2.
+Attention Bias and Constraint Drift.
+MACI avoids relying on a single LLM to execute complex, multi-step reasoning sequentially.
+Instead, it utilizes small collaborative agents that enjoy two key benefits: independence and well-defined input/output protocols (ensuring specificity and quality) for specific tasks.
+These agents operate within restricted context windows of e.g., 1k tokens, which physically limits attention bias and ensures that earlier constraints are not overridden by recent context.
+By logically scoping problems and physically constraining context, MACI preserves global feasibility and mitigates cognitive tunneling.
+Lack of Common Sense Integration.
+MACI incorporates a Common Sense Integration Agent and other specialized agents to identify implicit constraints and augment plans with practical, domain-specific knowledge.
+This ensures that generated plans are realistic, comprehensive, and aligned with real-world conditions.
+Through its innovative architecture, MACI overcomes the inherent limitations of LLMs, enhancing their capacity for deliberate reasoning and planning.
+In subsequent sections, we demonstrate MACI’s effectiveness through evaluations in complex scenarios, such as the Traveling Salesman Problem (TSP) and a multi-layered dinner planning task.
 2 Related Work
@@
 Gödel’s second incompleteness theorem (Kennedy, 2008; Gödel, 1967) established that no consistent formal system can prove its own consistency.
-We extend this insight to LLMs, which operate on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation.
-Table 1 illustrates how MACI’s distributed validation architecture addresses these fundamental limitations.
-Current multi-agent systems (MAS) primarily serve as integration platforms for coordinating multiple LLMs.
+This principle extends to LLMs, which rely on probabilistic rather than axiomatic foundations, making them inherently incapable of reliable self-validation.
+To address this, MACI employs a distributed validation architecture, where independent agents validate externally the output, bypassing the self-referential loops that may lead to inconsistencies.
+In formal systems, consistency proofs require a “higher-order” system.
+Analogously, MACI provides a validation framework that operates as a higher-order metasystem for LLM output.
+By decoupling planning from validation, MACI mirrors the separation needed in formal systems, where validation is performed independently to avoid conflicts and errors.
+Moreover, MACI advances the state of the art in multi-agent systems by addressing challenges that existing frameworks have not fully resolved.
+Current multi-agent systems (MAS) primarily function as integration platforms for coordinating multiple LLMs.
 Notable frameworks include Microsoft’s AutoGen (Wu et al., 2024), the Multi-LLM Agent Debate Framework (Du et al., 2023; Chang, 2023, 2024a, 2024b), LangGraph and CrewAI (LangChain AI, 2024; Moura, 2024), XAgent (Xia et al., 2023), and CAMEL (Li et al., 2023).
-Although these frameworks excel at agent coordination, they focus on task distribution rather than comprehensive constraint management required for complex planning.
-MACI’s meta-planning module (Section 4.1) addresses these limitations by acting as a metasystem that coordinates specialized agents.
-By decoupling planning and validation functions, MACI creates a validation framework analogous to the higher-order system required to demonstrate formal consistency while maintaining awareness of interdependent constraints.
-3 Case Study: Can LLMs Handle Planning
-This section begins by presenting various types of planning, including strategic approaches and tactical methodologies.
-We then present experimental results on a scheduling problem to illustrate how current LLMs fall short.
-Based on our diagnoses, we propose remedies in Section 4.
-Planning Types
-Planning can be classified into seven types. The first three are considered more strategic and the rest are tactic methodologies.
-  1. 1.
-    Sequential planning: Creates time-ordered schedules meeting goals/constraints via formal temporal representations (Allen & Hayes, 1989)
-  2. 2.
-    Proactive planning: Anticipates future scenarios to preemptively address issues through deliberate foresight (Cox & Veloso, 1998)
-  3. 3.
-    Retrospective planning: Improves future plans by analyzing past experiences via case-based reasoning (Kolodner, 1993)
-  4. 4.
-    Adaptive planning: Adjusts plans to changing conditions while preserving core constraints (Hammond, 1990)
-  5. 5.
-    Reactive planning: Prioritizes immediate actions over long-term goals in dynamic, uncertain environments (Georgeff & Lansky, 1987)
-  6. 6.
-    Predictive planning: Uses data-driven models to forecast scenarios and optimize decisions (Kushmerick et al., 1995)
+While these frameworks excel in agent coordination, they prioritize task distribution over the comprehensive constraint management necessary for complex planning.
+MACI bridges this gap by integrating a meta-planning module with independent agents that validate constraints, enabling robust and adaptable solutions in dynamic real-world scenarios.
+The meta-planner constructs task-specific dependency graphs that encode inter-agent constraints, ensuring precise role allocation while maintaining global feasibility.
+Meanwhile, validation agents, operating independently of the planning process, monitor for errors and inconsistencies stemming from probabilistic output, ensuring alignment with task objectives.
+This separation of roles mitigates cognitive tunneling and enhances adaptability, allowing MACI to dynamically respond to real-time disruptions such as resource shortages or environmental changes.
+By integrating these advanced mechanisms, MACI goes beyond existing MAS frameworks to provide a cohesive architecture for complex reasoning and planning.
+It ensures a high degree of scalability and robustness, making it suitable for applications ranging from logistical optimization to adaptive decision-making in uncertain environments.
+3 Case Study: Illuminating LLM Limitations
+Planning methodologies can be broadly categorized into sequential and reactive approaches.
+Sequential planning involves creating time-ordered schedules (Allen & Hayes, 1989), anticipating future scenarios (Cox & Veloso, 1998), and leveraging past experiences for improvement (Kolodner, 1993).
+Reactive planning, on the other hand, focuses on adapting to changing conditions (Hammond, 1990), prioritizing immediate actions in dynamic environments (Georgeff & Lansky, 1987), and using data-driven models to forecast scenarios (Kushmerick et al., 1995).
+This section highlights the limitations of current LLMs in planning tasks through experiments in two contexts: a scheduling problem illustrating shortcomings in sequential planning, and a dynamic resource allocation problem revealing challenges in reactive planning.
+Based on these observations, we propose remedies in Section 4.
 Problem Statement
@@
     When should Grandma be picked up by who?
-Using this problem, three key types of planning conducted by GPT4o and DeepSeek are examined: sequential, proactive/reactive, and retrospective.
+Using this problem, we examine sequential and reactive planning performed by GPT-4o and DeepSeek.
 Table 1: DeepSeek Proposed Schedule
@@
 2:30pm Emily lands at BOS. Emily
-3:00pm James arrives at airport to pick up Emily J
+3:00pm X James arrives at airport to pick up Emily J
 3:00pm Michael arrives at home and leaves to pick up Grandma. Michael
-3:15pm Michael at Grandma’s home. M
+3:15pm X Michael at Grandma’s home. M
 3:45pm Grandma arrives home; Michael returns. M
@@
 5:00pm Michael leaves to pick up Grandma M
-5:30pm Michael arrives home with Grandma M
+5:30pm X Michael arrives home with Grandma M
 6:00pm Dinner is served All
@@
 “someone”), creating brittleness under reduced slack time (e.g., delays).
-Analysis (Appendix A) links both errors to flawed reasoning in constraint interpretation.
+Analysis (with detailed execution in Appendix A) links both errors to flawed reasoning in constraint interpretation.
 Diagnoses: Common-Sense Constraints and Isolated Processing Syndrome
@@
 This syndrome manifests itself in two critical ways: the system either misses obvious optimizations or proposes solutions that violate the stated constraints, leading to an infeasible or suboptimal plan.
-3.2 Study #2 Proactive and Reactive Planning
+3.2 Study #2 Reactive Planning
 Real-world scenarios do not always follow plans precisely.
@@
 Michael departs to pick up Emily from airport Michael
-3:00pm Sarah departs to pick up Grandma Sarah
+3:00pm X Sarah departs to pick up Grandma Sarah
 3:30pm Arrive at Grandma’s house Sarah
@@
 More specifically, GPT-4o assigned Sarah to monitor the oven, but missed reallocating this task to Grandma earlier, preventing Sarah from serving as an additional driver, a missed efficiency gain.
-3.3 Study #3 Retrospective Planning
-Retrospective planning improves future strategies by learning from past decisions.
-Both DeepSeek and GPT-4o lacked flexibility even after error correction: DeepSeek relied on rigid templates, and GPT-4o’s initial assignment (Sarah homebound in Table 3) limited later options.
-Missed Opportunities During James’ 3-Hour Delay:
-  • Sarah could have picked up Grandma earlier, freeing her for contingencies like Emily’s transport.
-  • Michael could have left NYC at 10:00 AM to reach Emily by 2:30 PM.
-In Section 4, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
 3.3 Summary of LLM Limitations in Planning
@@
 LLMs miss implicit real-world knowledge (luggage claim times, rental car logistics) requiring explicit specification of human-obvious constraints (airport-terminal proximity).
+In Sections 4 and 5, we show how MACI’s meta-planner effectively revises plans by debugging errors and adapting to dynamic constraints.
 4 MACI Framework Specification
-MACI implements a three-layer architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning.
-The top two layers represent key innovations, while the bottom layer leverages existing multi-agent system (MAS) infrastructure.
-4.1 Three-Layer Architecture
-Meta-Planner: The Planning Generator
+MACI implements a three-component architecture to address current LLM limitations: metacognitive constraints, attention bias, and gaps in common-sense reasoning.
+Each component plays a distinct role in enabling robust and adaptable planning capabilities.
+4.1 Three-Component Architecture
+Meta-Planner Component
 The meta-planner $\mathcal{MP}$ functions as a higher-order planner that generates task-specific planning systems:
-$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},$
+$\mathcal{MP}:(\mathcal{O},\mathcal{C_{E}})\rightarrow\mathbf{W},\vspace{-.05in}$
 where $\mathbf{W}$ represents a planning system composed of specialized, coordinated agents.
-Similar to how a compiler generator produces compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements.
+Similar to a compiler generator producing compilers from specifications, $\mathcal{MP}$ constructs agent networks from task requirements.
 It analyzes objectives, identifies required roles and dependencies, selects appropriate agents, and establishes interaction protocols.
-This produces a complete workflow template defining both the planning state space and the agent coordination mechanisms needed to solve the given task.
-Agent Repository Layer
-This layer maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface.
-By dividing cognitive tasks among agents, this architecture ensures a complete knowledge of the constraints without overwhelming individual components.
-System Infrastructure Layer
-Built on open-source MAS frameworks, the infrastructure layer supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling.
-4.2 Agent Repository Layer Design
-4.2.1 Agent Registration and Search Protocol
-Agents register in the repository with five attributes: input/output protocol ($P$), type ($t$), capability vector ($c$), context window size ($w\leq 1k$ tokens) and user rating ($r$).
-These attributes enable $\mathcal{MP}$ to match task requirements to agent capabilities efficiently.
-$\mathcal{MP}$ searches for agents by: 1) matching task requirements to capability descriptions ($c$), 2) filtering by agent type ($t$) and user ratings ($r$), and 3) verifying protocol compatibility ($P$).
+This produces a workflow template that defines the planning state space and the coordination mechanisms needed to solve the task.
+Agent Repository Component
+This component maintains a distributed collection of planning agents, each designed with a restricted context window and specialized interface.
+By dividing cognitive tasks among agents, the repository ensures a complete representation of constraints without overwhelming individual components.
+The meta-planner queries this repository to select agents for specific roles and dependencies based on task requirements.
+System Infrastructure Component
+Built on open-source multi-agent system (MAS) frameworks, the infrastructure component supports essential operations such as agent registration, message routing, resource allocation, and deployment scaling.
+This foundation provides the necessary runtime environment for executing and monitoring the generated workflows.
+4.2 Agent Repository Design
+The agent repository in MACI serves as a structured database, enabling efficient registration, retrieval, and matching of agents to task requirements.
+By categorizing agents into common agents and specialized agents, the repository supports both generalized functionality and domain-specific expertise, as outlined in Section 4.1.
+4.2.1 Lightweight, Independent Agent Design
+MACI avoids relying on a single LLM to execute complex, multi-step reasoning sequentially.
+Instead, it utilizes small, independent agents that adhere to strict efficiency and modularity principles.
+These agents operate with well-defined input/output protocols and are constrained to restricted context windows to mitigate attention bias and prevent earlier constraints from being overridden by recent context.
+By scoping problems logically and constraining context physically, MACI ensures that each agent processes only the task-relevant information needed for its specific role.
+This design prevents cognitive tunneling, maintains global feasibility, and enhances robustness in dynamic environments.
+4.2.2 Agent Registration and Specifications
+Each agent is registered in the repository using a standardized protocol buffer that encodes the following attributes:
+  • Input/output protocol ($P$): Defines the data format and expected interactions for seamless communication.
+  • Agent type ($t$): Specifies whether the agent is common or specialized.
+  • Capability vector ($c$): Encodes the agent’s functional capabilities, constraints, and operating conditions.
+  • Context window size ($w$): Ensures that each agent operates within a restricted buffer ($w\leq 1k$ tokens) to prevent attention bias and excessive information retention.
+  • Computational efficiency constraint ($e$): Agents are lightweight, avoiding unnecessary memory usage or processing delays.
+  • User rating ($r$): Tracks historical performance evaluations to prioritize reliable agents during selection.
+The meta-planner retrieves agents from the repository using a three-step matching process:
+  1. 1.
+    Task-to-capability matching: Filters agents based on their capability vector ($c$) and task-specific requirements.
+  2. 2.
+    Protocol verification: Ensures compatibility of input/output protocols ($P$) between selected agents to prevent communication errors.
+  3. 3.
+    Agent ranking: Ranks agents by their relevance, efficiency, and historical user rating ($r$) to select the optimal candidates.
+This structured retrieval mechanism ensures that MACI efficiently scales to complex planning problems without requiring predefined agent hierarchies.
+By leveraging protocol buffers and a structured repository, MACI achieves both modularity and adaptability, allowing new agents to be introduced seamlessly while maintaining coherence across multi-agent interactions.
 4.2.3 State Space and Agent Design
-General State Space
-Table 5: State Space Dimensions and Components
-Dimension Core Components Example States
-Who
-• Actor ID
-• Current Roles
-• Role Constraints
-• Role Changes
-• Driver vs Passenger
-• Supervisor vs Worker
-• Buyer vs Seller
-• Multiple role conflicts
-Where
-• Current Position
-• Target Location
-• Transition Points
-• Access Rules
-• Physical locations
-• Virtual positions
-• State transitions
-• Boundary constraints
-When
-• Time Points
-• Duration
-• Deadlines
-• Probability
-• Event timestamps
-• Process duration
-• Completion times
-• Delay likelihood
-What
-• Methods
-• Requirements
-• Constraints
-• Costs
-• Tools/Vehicles
-• Tickets/Permits
-• Capacity limits
-• Time-money trade-offs
-Why
-• Rationale
-• Dependencies
-• Risks
-• Alternatives
-• Decision basis
-• Causal chains
-• Failure modes
-• Backup plans
-Table 5 presents five dimensions of state space, Who, Where, When, What, and Why, that form the basis of planning tasks.
-These dimensions are general enough to apply across domains, with travel planning serving as an illustrative example.
+Tasks in MACI are modeled in a general five-dimensional state space to ensure comprehensive representation of constraints and dependencies. These dimensions include:
+  1. 1.
+    Who (Actors): Identifies roles, constraints, and transitions between agents or individuals.
+  2. 2.
+    Where (Location): Tracks physical or logical positions, transitions, and access rules.
+  3. 3.
+    When (Time): Captures temporal constraints such as deadlines, durations, and time points.
+  4. 4.
+    What (Resources): Manages resource availability, constraints, and associated costs.
+  5. 5.
+    Why (Logic): Encodes rationale, dependencies, and risk assessments for decision-making.
+This structured state space allows the meta-planner to generate workflows that account for all relevant constraints and dependencies across diverse domains.
+4.2.4 Agent Roles in State Space Management
 Common Agents
-Common agents are designed to handle core planning and validation tasks that are universally applicable across domains.
-These agents provide foundational capabilities that ensure the consistency, feasibility, and robustness of planning systems, regardless of the specific application.
-Their roles align with the general state space dimensions of Who, Where, When, What, and Why, ensuring that all essential aspects of planning are addressed.
-For a list of common agents, please refer to Appendix D.2.
-These agents work collaboratively to ensure robust planning and validation across diverse domains.
-Although the list of common agents is not exhaustive, it provides a strong foundation to address most planning tasks.
-New common agents can be introduced as the need arises, provided they offer capabilities that can be reused across multiple domains.
-For example, an Adaptation Agent could be added to dynamically adjust constraints in response to evolving task environments.
-Specialized Agents
-Specialized agents extend the capabilities of MACI by addressing domain-specific requirements that are not generalizable across all tasks.
-These agents are designed to implement detailed knowledge and expertise tailored to specific applications, ensuring that MACI remains adaptable and effective for diverse planning challenges.
-Specialized agents typically focus on tasks requiring unique constraint handling, optimization, or analysis.
-Examples of specialized agents include the following.
-  • Travel Planning Agent: Optimizes itineraries, manages transportation options, and ensures seamless transitions between locations.
-  • Financial Risk Agent: Monitors market trends, assesses risks, and validates investment strategies.
-  • Healthcare Resource Agent: Allocates medical supplies, schedules staff, and optimizes patient care pathways.
-  • Supply Chain Agent: Manages logistics, tracks inventory, and optimizes distribution routes.
-Specialized agents adhere to architectural principles that maintain consistency and integration within MACI:
-  • Restricted Context Windows: Each agent operates within a limited context to ensure efficiency and focus ($w\leq 1k$ tokens).
-  • Protocol Buffer Interfaces: Standardized interfaces enable seamless communication between agents.
-  • Coordination with Common Agents: Specialized agents leverage the capabilities of common agents for validation, compliance, and constraint management.
-  • Domain-Specific Constraints: Specialized agents are equipped to handle unique requirements and challenges specific to their domain.
-4.2.3 Agent Coordination and State & Constraint Management
-MACI ensures seamless coordination between agents across state dimensions.
-For example, in travel planning, when a Schedule Agent detects a flight connection constraint, it triggers updates to Location, Resource, and Role Manager Agents.
-These updates propagate through a state transition protocol, where the proposed state change is validated against global and local constraints:
-$T:S_{t}\rightarrow S_{t+1}\text{ subject to }C_{\text{global}}\land C_{\text{local}}$
-4.2.4 Validation Protocols
-The validation agent ensures state consistency by decomposing global constraints into dimension-specific checks.
-For example, when booking a flight, the temporal agent validates the temporal constraint, the spatial feasibility is checked by the location agent, and the availability of seats is validated by the resource agent.
-Common sense agents identify overlooked constraints, such as luggage claiming time, inter-terminal walking time, and rental car pickup time.
-If validation fails, the system triggers a recovery process to generate alternative solutions.
-Validation protocols can be updated when unexpected events occur to improve retrospective planning.
-4.2.5 Other Considerations
-Due to space limitation, we document in Appendix C additional considerations including cross-domain generalization, enhancement and integration of common sense knowledge, scalability, and a list of future enhancements when more applications will be supported.
+Common agents are designed to handle foundational planning tasks that align with MACI’s state space dimensions (Who, Where, When, What, Why).
+These agents provide general-purpose functionality that ensures consistency, feasibility, and robustness across planning tasks.
+Their primary responsibilities include:
+  • Constraint Validation Agents: Ensure adherence to temporal, spatial, and resource constraints by verifying the feasibility of generated plans.
+  • Common Sense Integration Agents: Identify implicit constraints that may be overlooked, such as transition times, dependencies, or practical limitations.
+  • Adaptation Agents: Dynamically adjust plans in response to changes in task environments, such as resource delays or evolving requirements.
+  • Performance Evaluation Agents: Assess the quality and efficiency of proposed plans relative to predefined metrics, ensuring continuous improvement.
+By addressing these tasks, common agents form the backbone of MACI’s planning architecture.
+Their modular design enables reuse across multiple domains, and their collaborative functionality ensures they work seamlessly with specialized agents to maintain global consistency and coherence within the planning workflow.
+Task-Specific Agents
+These agents cater to domain-specific requirements, including task-dependent data and knowledge augmentation, selecting and optimizing planning algorithms, safety and ethics assessment, and emergency response optimization.
+By leveraging domain expertise, specialized agents extend the capabilities of common agents, enabling MACI to address specialized planning challenges with precision and adaptability.
+4.2.5 Seamless Integration and Scalability
+The repository’s standardized agent specifications and matching mechanism enable MACI to scale efficiently across domains.
+By leveraging modular designs and protocol buffers, new agents can be integrated seamlessly into existing workflows, ensuring adaptability and extensibility without compromising performance or consistency.
 4.3 Meta-Planner: Planning a Planner to Plan
-The meta-planner $\mathcal{MP}$ operates as a higher-level planner that constructs other planners, which are actionable workflows for specific tasks.
+The mission of the meta-planner $\mathcal{MP}$ is to construct a planner that generates an actionable workflow for a given task.
+It does so by analyzing task objectives, identifying roles and constraints, and organizing agents into a structured execution plan.
+This three-phase approach ensures that every agent and dependency is optimally placed, refined, and validated, leading to robust, task-specific workflows.
 4.3.1 The Meta-Planner Algorithm
-$\mathcal{MP}$ is specified in Algorithm #1. A list of all symbols used in the specifications is provided in Table 20 in Appendix D.
-A workflow network $\mathbf{W}=(\mathcal{N},\mathcal{E})$, represented as a directed graph, consists of roles as nodes and their dependencies as edges.
-The nodes $n\in\mathcal{N}$ denote the roles performed by human actors (e.g.
-driver, cook, supervisor), while the edges $e_{ij}\in\mathcal{E}$ between the nodes $n_{i}$ and $n_{j}$ capture the monitored dependencies (e.g., temporal, spatial constraints, data dependencies and safety requirements for supervision).
-$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n}\text{and }\mathcal{E}=A^{*}_{e}.\vspace{-.1in}$ (1)
+The meta-planner operates as a higher-order planning system that formulates workflows as directed graphs:
+$\mathbf{W}=(\mathcal{N},\mathcal{E}),\text{ where }\mathcal{N}=A^{*}_{n},\mathcal{E}=A^{*}_{e}.$ (1)
+Here, $\mathcal{N}$ denotes roles assigned to agents, and $\mathcal{E}$ represents dependencies between roles, including constraints such as timing, data flow, and supervision requirements.
 Algorithm 1 $\mathcal{MP}$: Planner for Planning a Plan
@@
   // Phase 1: Network Construction
-  1. $\mathcal{N}$: Extract roles & capabilities from $\mathcal{O}$ (Eq. 2)
-  2. $\mathcal{E}$: Identify node dependencies from $\mathcal{C_{E}}$ (Eq. 3)
+  1. Extract roles $\mathcal{N}$ from $\mathcal{O}$ (Eq. 2)
+  2. Identify dependencies $\mathcal{E}$ from $\mathcal{C_{E}}$ (Eq. 3)
   // Phase 2: Agent Assignment
-  3. Assign role agents: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
-  4. Assign edge agents: $\forall e_{ij}\in\mathcal{E}$, $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
+  3. Assign agents to nodes: $\forall n\in\mathcal{N}$, select $\alpha_{n}\in\mathbf{A_{n}}$ (Eq. 4)
+  4. Assign agents to dependencies: $\forall e_{ij}\in\mathcal{E}$, select $\alpha_{ij}\in\mathbf{A_{e}}$ (Eq. 5)
   // Phase 3: Iterative Refinement
   while improvement in $V(\mathbf{W},\mathcal{M})$ do
-     // Role Assignment Optimization
      for all $n\in\mathcal{N}$ do
-        Update role-person mappings via $f_{\text{role}}(n,\mathcal{P})$
+        Update role-person mappings $f_{\text{role}}(n,\mathcal{P})$
      end for
-     // Dependency Monitoring
      for all $e\in\mathcal{E}$ do
@@
      end for
-     // Validation
      if $V(\mathbf{W_{new}},\mathcal{M})>V(\mathbf{W_{current}},\mathcal{M})$ then
@@
   end while
-  return $\mathbf{W^{*}}=\mathbf{W{current}}$
-$\mathcal{MP}$ constructs a planner in three phases. First, it builds the network structure by identifying roles and their dependencies, forming the foundation of the workflow.
-Second, it assigns appropriate agents to manage both roles and dependencies.
-Each node $n\in\mathcal{N}$ is managed by a designated role agent $\alpha_{n}$.
-Similarly, each edge $e_{ij}\in\mathcal{E}$, representing a dependency between nodes, is managed by an agent $\alpha_{ij}$, which connects the node agents $\alpha_{i}$ and $\alpha_{j}$.
-These edges are enriched with metadata that may include details such as data types, communication protocols, timing constraints, and security requirements.
-Finally, the generated planner iteratively refines the workflow.
-This process involves optimizing role assignments and continuously monitoring dependencies until the validation score stabilizes, ensuring a well-validated workflow.
-4.3.2 Meta-Planning Components
-Let us clearly distinguish between roles (requiring human qualifications) and monitoring tasks (requiring capabilities).
+  return $\mathbf{W^{*}}=\mathbf{W_{current}}$
+4.3.2 Meta-Planning Design Elements
 Role and Qualification Analysis
-$\mathcal{MP}$ analyzes objectives $\mathcal{O}$ to identify required roles and their qualifications:
+The meta-planner extracts roles from task objectives and maps them to required qualifications:
 $\text{map}_{\text{role}}:\mathcal{O}\rightarrow\{(n_{i},q_{i})\}$ (2)
-where $n_{i}$ represents a role and $q_{i}$ its required qualifications. For example, $(n_{\text{driver}},q_{\text{license}})$ or $(n_{\text{cook}},q_{\text{experience}})$.
+where $n_{i}$ represents a role and $q_{i}$ its required qualifications (e.g., a driver requires a license, a cook requires experience).
 Constraint Management
-The framework maintains a global constraint set $C$ composed of:
+Constraints govern role interactions and dependencies. The framework maintains a global constraint set:
 $C=C_{E}\cup C_{I}\cup C_{D}$ (3)
-where $C_{E}$ represents explicit constraints from problem specification, $C_{I}$ denotes implicit constraints identified by Common Sense Agent, and $C_{D}$ represents derived constraints from agent interactions.
+where $C_{E}$ represents explicit constraints from task specifications, $C_{I}$ denotes implicit constraints identified by common sense agents, and $C_{D}$ represents derived constraints from agent interactions.
 Agent Assignment
-Two types of agents, node and edge, are selected from the repository $\mathbf{A}$:
-1. Node Monitoring Agents:
+Two categories of agents are assigned based on task requirements:
+  • Node Agents (Role Execution):
     $A^{*}_{n}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{n_{j}}\text{dist}(q_{j},A_{i}.\text{capabilities})$ (4)
-These agents verify role qualifications and manage people-role assignments and transitions.
-2. Edge Monitoring Agents:
+    These agents are responsible for fulfilling role qualifications and managing people-role assignments.
+  • Edge Agents (Dependency Management):
     $A^{*}_{e}=\operatorname*{arg\,min}_{A_{i}\in\mathbf{A}}\sum_{e_{j}}\text{dist}(c_{j},A_{i}.\text{capabilities})$ (5)
-These agents monitor dependencies between roles, where $c_{j}$ represents constraint types (temporal, spatial, etc.).
-$\mathcal{MP}$ iteratively refines the network by validating workflow constraints and performance and replacing or adding agents, i.e., checking and adding nodes and edges.
+    These agents ensure dependencies between roles are correctly maintained, such as time constraints, spatial relations, and supervisory requirements.
 4.4 Workflow Execution Framework
-The workflow $\mathbf{W^{*}}$ requires a runtime system for execution.
-In this work, we evaluate $\mathbf{W^{*}}$ by entering $\mathbf{W^{*}}$ along with the problem statement into an LLM (DeepSeek and GPT4o).
-A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ must be performed manually.
-5 Evaluating $\mathcal{MP}$’s Planning Capabilities
+The final workflow $\mathbf{W^{*}}$ must be executed in a runtime environment.
+In this work, we evaluate $\mathbf{W^{*}}$ by entering it into an LLM (e.g., GPT4o) alongside the problem statement.
+A key limitation is that the feedback loop for refining $\mathbf{W^{*}}$ is currently manual, requiring iterative adjustments to optimize execution.
+Future research will focus on automating this process to enhance adaptability and efficiency.
+5 Evaluating $\mathcal{MP}$ vs. Independent LLMs
 To assess $\mathcal{MP}$’s performance and adaptability, we adopted a dual-approach experimental structure.
@@
 The second involves the Thanksgiving Dinner Planning problem, showcasing $\mathcal{MP}$’s ability to handle complex, real-world challenges with cross-thread dependencies and dynamic adaptability.
-Traveling Salesperson Problem (TSP)
-The TSP experiment benchmarks $\mathcal{MP}$ against planners (Claude, DeepSeek, GPT4o) and their $\mathcal{MP}$ integrations.
-Metrics include: 1) solving quality, and 2) solution optimality.
-Detailed results are in Appendix E.
-The conclusion is that though the task is simple, $\mathcal{MP}$ helps improve results for monitoring execution, validating results, and suggesting algorithmic approaches.
-Thanksgiving Dinner Planning
-This problem, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraints and dependencies.
-It involves multiple agents performing interdependent tasks, representing a complex multi-agent scenario.
-Planning performance is compared across three configurations: DeepSeek + $\mathcal{MP}$, GPT4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$. Metrics include:
+Due to space constraints, detailed results for these experiments are provided in Appendices D and E, respectively.
+5.1 Traveling Salesperson Problem (TSP)
+The TSP experiment benchmarks $\mathcal{MP}$ against standalone planners (Claude, DeepSeek, GPT-4o) and their $\mathcal{MP}$-integrated counterparts.
+The metrics include solution quality and optimality.
+Result Summary
+Without $\mathcal{MP}$, DeepSeek performs best, while Claude and GPT-4o struggle, each exceeding the optimal travel time by more than $10\%$.
+With $\mathcal{MP}$, Claude requires two iterations to reach the optimal distance, while both GPT-4o and DeepSeek solve the problem in a single attempt.
+Although TSP involves a straightforward, single-thread planning process, $\mathcal{MP}$ still provides notable enhancements. Again, see Appendix D for details.
+5.2 Thanksgiving Dinner Planning
+This task, detailed in Section 3, evaluates $\mathcal{MP}$’s ability to generate workflows $\mathbf{W^{*}}$ with enhanced constraint and dependency management in the MACI setting.
+Unlike TSP, this problem involves multiple interdependent agents, introducing complex coordination challenges.
+Planning performance is assessed across three configurations: DeepSeek + $\mathcal{MP}$, GPT-4o + $\mathcal{MP}$, and Claude + $\mathcal{MP}$.
+The prior results in Section 3 show that all LLMs fail the task when executed independently.
+Evaluation metrics include:
 […]
 where flexibility measures slack time incorporated to handle unexpected events.
-5.1 Meta-Planning for Thanksgiving Event
+5.2.1 Meta-Planning for Thanksgiving Event
 Following Algorithm 1, $\mathcal{MP}$ generates workflows with:
@@
   • Implicit constraint edges from common-sense analysis.
-The planner monitors nodes and edges, enabling dynamic adjustments. Full specifications are in Appendix A.
+The planner monitors nodes and edges, enabling dynamic adjustments. The full specifications are in Appendix E.
 Evaluation Scenarios
@@
   • Incorporating common-sense constraints (e.g., fatigue, social preferences),
-In reactive planning, $\mathcal{MP}$ integrates an early information agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
-5.2 Experimental Results
+In reactive planning, $\mathcal{MP}$ integrates an alert agent to detect flight delays at departure, enabling timely workflow updates and demonstrating adaptability.
+5.2.2 Experimental Results
 Sequential Planning Performance
 With $\mathcal{MP}$’s enhanced workflow $\mathbf{W^{*}}$, all three LLMs successfully generated feasible solutions, a significant improvement over their previous failures with the original problem specification.
-As illustrated in Table 6, DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations.
+Table 4 summarizes the detailed schedules documented in Tables 12, 13, and 14, in Appendix E.8.
+DeepSeek demonstrated superior scheduling efficiency by optimizing James’s airport wait time for Emily’s pickup, requiring only two iterations.
 While GPT4o eventually produced a valid solution in three iterations, it created suboptimal travel patterns by having Michael make separate trips.
@@
 Reactive Planning Performance
-GPT4o failed to produce a valid solution, violating three critical constraints without the ability to recognize these violations.
-Reactive Planning Performance
 The flight delay scenario revealed significant differences between LLMs’ capabilities.
 DeepSeek demonstrated superior spatial reasoning by routing Michael directly to the airport, an insight that should have come from $\mathcal{MP}$’s common-sense spatial reasoning.
-This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix A.9), which DeepSeek leveraged for immediate replanning.
+This unexpected ability to improve workflow highlights the synergy between $\mathcal{MP}$ and LLM —$\mathcal{MP}$ provided early alert through its information agent (Table 15 in Appendix E.9).
+Table 5 summarizes the detailed schedules documented in Tables 16, 18, and 19, in Appendix E.9.
+DeepSeek leveraged the early alert at 10:00 AM for immediate replanning.
 In contrast, Claude produced two feasible plans but missed the 10:00 AM alert in $\mathbf{W^{R}}$, starting its schedule at 1:00 PM and missing opportunities for proactive actions like early Grandma pickup to free Sarah as a driver.
-GPT4o failed entirely, producing three constraint violations it couldn’t recognize, preventing further improvements.
-Table 7: Reactive Planning Performance (Flight Delay)
+GPT4o failed entirely, producing three constraint violations it could not recognize, preventing further improvements.
+Table 5: Reactive Planning Performance (Alert: flight delay)
 LLM # Notable Features
 DeepSeek 3 Smart routing of Michael directly to airport; efficient travel patterns
-GPT4o Failed Failed to maintain critical constraints; unable to recover
+GPT4o X Failed to maintain critical constraints; unable to recover
 Claude 3 Two valid plans with different trade-offs; longer wait times
-6 Conclusions with Impact Statements
-Current artificial intelligence systems excel at pattern matching but struggle with deliberate analytical reasoning required for complex planning.
-While Large Language Models (LLMs) demonstrate impressive capabilities, they face three critical limitations: inability to validate their own outputs (paralleling Gödel’s incompleteness theorems), inconsistent maintenance of constraints during planning, and inadequate integration of common-sense knowledge.
-Our experiments demonstrate that existing multi-agent systems, focused primarily on coordination, fail to address these limitations.
-MACI (Multi-Agent Collaborative Intelligence) advances the state of the art through three key contributions:
+6 Conclusion with Impact Statements
+This research introduces Multi-Agent Collaborative Intelligence (MACI), a framework designed to overcome fundamental limitations in current LLMs for reasoning and planning.
+By reimagining computational intelligence through a hierarchical, distributed architecture, MACI represents a structured approach to complex problem-solving.
+To validate MACI’s capabilities, two experiments tested its effectiveness in adaptive planning and constraint management.
+The first experiment, based on the Traveling Salesperson Problem (TSP), demonstrated MACI’s ability to generate globally optimized workflows that guide LLMs in solving this classical problem.
+The second, a Thanksgiving Dinner Planning task, highlighted MACI’s ability to resolve intricate cross-thread dependencies and dynamically adjust to evolving constraints.
+MACI introduces three fundamental innovations that redefine computational reasoning:
+Structured Meta-Planning for Constraint-Aware Execution:
+The framework implements a structured meta-planner that explicitly separates planning from execution.
+This architectural approach dismantles the attention-driven limitations that have historically constrained LLM performance, enabling a more deliberate and controlled planning mechanism.
+Distributed Validation for Reliable Decision-Making:
+MACI establishes a distributed validation mechanism that enhances system reliability.
+By incorporating independent verification agents, the framework transcends the probabilistic limitations of traditional language models, introducing a new paradigm of self-checking computational intelligence.
+Proactive Multi-Agent Coordination for Adaptive Planning:
+The system enables a proactive approach to multi-agent coordination.
+Unlike previous static collaboration frameworks, MACI facilitates real-time constraint resolution and dynamic replanning, significantly advancing adaptive computational reasoning.
+By redefining intelligent planning, MACI lays the foundation for AI systems that operate with greater precision, adaptability, and contextual awareness, bridging the gap between static LLM reasoning and dynamic real-world decision-making.
+Appendices
+Appendix A Validation and Recovery Protocols
+The validation protocol implements a multi-stage process for ensuring state consistency. When any agent proposes a state change, the validation agent initiates a sequence of checks:
+$\text{validate}(s_{t}\rightarrow s_{t+1})=\begin{cases}\text{true}&\text{if all checks pass}\\ \text{false}&\text{if any check fails}\end{cases}\vspace{-.1in}$ (6)
+The protocol begins with pre-validation.
+Before a state transition starts, the validation agent queries relevant agents about preconditions.
+For a travel booking, temporal agent verifies the proposed times fit within existing schedules.
+Spatial agent confirms the physical feasibility of movements between locations.
+Role agent checks if all actors can perform their assigned functions.
+During the transition, the protocol maintains atomic operations.
+The validation agent tracks changes across all state dimensions, ensuring partial updates cannot create inconsistent states.
+If the temporal agent approves a flight time but the resource agent finds insufficient seats, the entire transition fails and rolls back.
+Post-validation examines the resulting state.
+The validation agent verifies that all constraints remain satisfied after the change.
+Common sense agent reviews the new state for practical issues that formal checks might miss.
+Strategy agent confirms the transition aligns with overall planning objectives.
+When validation fails, the protocol triggers a structured recovery process:
+$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (7)
+Recovery begins by logging the failure cause and violated constraints.
+The strategy agent then works with domain agents to generate alternative proposals that satisfy the constraints.
+This might involve relaxing non-critical constraints or exploring different approaches to meet the planning objectives.
+A.1 Operations Research Techniques in Validation Protocols
+The validation protocols described above align closely with established methods in operations research (OR). Some relevant techniques include:
+  • Constraint Programming (CP): Focuses on solving combinatorial problems by enforcing constraints, ensuring consistency across dimensions such as temporal, spatial, and resource availability (Rossi et al., 2006).
+  • Mixed-Integer Linear Programming (MILP): Optimizes decision variables subject to linear constraints and objective functions, often used in scheduling and resource allocation (Wolsey & Nemhauser, 1998).
+  • Network Flow Algorithms: Validates feasibility and optimizes flows in networks by ensuring capacity, timing, and availability constraints are satisfied (Ahuja et al., 1993).
+  • Dynamic Programming (DP): Decomposes problems into sequential subproblems, useful for validating processes like inter-terminal walking or luggage claiming time (Bellman, 1957).
+  • Monte Carlo Simulation: Simulates scenarios to validate feasibility and robustness under uncertainty (Metropolis & Ulam, 1949).
+  • Robust Optimization: Focuses on solutions that remain feasible under uncertainty, ensuring plans adapt to disruptions (Ben-Tal & Nemirovski, 2009).
+Integration in Multi-Agent Systems
+The validation agent also employs techniques from multi-agent systems, such as:
+  • Blackboard Systems: A shared workspace for collaborative validation by different agents, ensuring global consistency (England & Engelmore, 1987).
+  • Consensus Protocols: Used for distributed validation, where agents negotiate to ensure all constraints are met (Ren & Beard, 2005).
+By combining these OR techniques with agent-based systems, the validation protocol ensures comprehensive and adaptive checks for workflow consistency.
+Future work can explore integrating heuristic methods, such as genetic algorithms or simulated annealing, to further enhance recovery processes.
+Appendix B $\mathsf{MACI}$ Additional Design Considerations
+B.1 Cross-Domain Generalization
+While the state space dimensions—Who, Where, When, What, and Why—are broad enough to cover diverse domains, additional customization may be required for unique applications.
+This section examines how MACI generalizes across domains like financial planning, healthcare logistics, and supply chain optimization.
+The travel planning example is illustrative, emphasizing how MACI dynamically adapts state spaces and agents to domain-specific requirements.
+B.2 Dynamic Agent Registration and Evolution
+This section explores how agents are dynamically developed, trained, and validated for new tasks.
+It discusses mechanisms for evaluating new agents and integrating them into the repository without retraining the entire system, ensuring scalability and adaptability.
+B.3 Scalability and Resource Efficiency
+As the number of agents and task complexity grows, MACI employs strategies to manage communication overhead and optimize agent interactions.
+This section details techniques for clustering agents and hierarchical coordination to maintain scalability.
+B.4 Empirical Evaluation Across Domains
+To demonstrate MACI’s adaptability, this section presents empirical results from applying the framework to multiple domains.
+Examples include financial portfolio management, urban traffic planning, and hospital resource allocation, highlighting MACI’s advantages over state-of-the-art systems.
+B.5 Challenges and Future Directions
+While MACI addresses many limitations of LLM-based planning, challenges remain in real-time coordination, implicit knowledge integration, and robust recovery mechanisms.
+This section proposes future research directions to enhance MACI’s performance and applicability to novel tasks.
+Appendix C Additional Tables and Figures
+C.1 Notation
+Table 6 presents all symbols used throughout this paper.
+Table 6: Symbol Definitions
+Symbol Definition Symbol Definition
+Basic Sets
+$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
+$\mathcal{C_{E}}$ Explicit constraints $\mathcal{C_{I}}$ Implicit constraints
+$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
+Workflow Components
+$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
+$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
+$n_{i}$ Individual role $e_{ij}$ Role dependency
+Agent Functions
+$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
+$V(\cdot)$ Validation function $\text{dist}(\cdot)$ Capability distance
+$\text{map}_{\text{role}}$ Role extraction $\text{map}_{\text{edge}}$ Dependency extraction
+Optimization
+$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
+$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
+C.2 Example Common Agents
   1. 1.
-    Meta-Planning Architecture: MACI generates specialized planning workflows by:
-    • Constructing role-node and dependency networks.
-    • Identifying implicit constraints through common-sense analysis.
-    • Establishing validation across distributed agents.
+    Role Manager Agent (Who): Tracks actors, their roles, and their associated constraints, ensuring that all role-based requirements are satisfied.
   2. 2.
-    Workflow Enhancement: As demonstrated in our experiments:
-    • Sequential Planning: Enhanced workflow $\mathbf{W^{*}}$ enabled all LLMs to generate feasible solutions.
-    • Reactive Planning: Early information agent in $\mathbf{W^{R}}$ enabled proactive replanning.
-    • Unexpected synergy: DeepSeek contributed spatial optimizations beyond MACI’s specifications.
+    Spatial Agent (Where): Manages geographic and location-based constraints, verifying transitions between physical or virtual locations.
   3. 3.
-    Empirical Validation: Our Thanksgiving dinner planning case study shows:
-    • Improving constraint satisfaction on LLMs.
-    • Effective handling of crises through early detection.
-    • Clear performance differences between LLMs in utilizing MACI’s capabilities.
-MACI demonstrates that addressing LLM limitations requires more than simple agent coordination.
-It demands a structured meta-planning architecture that enables comprehensive constraint awareness and validation.
-The framework’s success in complex temporal planning, coupled with unexpected improvements from LLM contributions, suggests a promising direction for developing more capable AI planning systems.
-Appendices
-Appendix A Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
+    Temporal Agent (When): Handles scheduling, timing, and deadlines, ensuring alignment with temporal constraints.
+  4. 4.
+    Resource Agent (What): Tracks real-world resources such as tools, vehicles, or financial instruments, managing capacity, availability, and associated costs.
+  5. 5.
+    Reasoning and Explanation Agent (Why): Maintains the rationale behind decisions, dependencies, and alternative plans, enabling consistent alignment with objectives and providing explanations for outcomes.
+  6. 6.
+    Common Sense Agent: Identifies implicit constraints, integrates practical knowledge, and ensures plans align with real-world considerations.
+  7. 7.
+    Constraint Validation Agent: Ensures that all constraints are satisfied and that proposed plans remain feasible.
+  8. 8.
+    Plan Evaluation Agent: Assesses the effectiveness of plans against predefined metrics and objectives.
+  9. 9.
+    What-If Testing Agent: Evaluates plan robustness by simulating alternative scenarios and analyzing their impact.
+  10. 10.
+    Compliance and Safety Agent: Monitors adherence to safety standards, ethical principles, and regulatory frameworks.
+Appendix D Traveling Salesman Problem Experiment
+D.1 General Problem Specification
+The TSP requires finding the shortest possible route visiting N locations exactly once, returning to the start:
+  • Inputs: N locations, distance matrix D[N][N]
+  • Output: Optimal tour T minimizing total distance
+  • Constraints: Each location visited once, return to start
+Computational Complexity - Brute Force
+For N locations:
+  • Number of possible tours = (N-1)!/2
+  • Time complexity = O(N!)
+  • Space complexity = O(N2)
+Solution Methods
+  1. 1.
+    Exact Methods: Representative methods are Branch and Bound (Land & Doig, 1960), Dynamic Programming (Bellman, 1962), and Integer Linear Programming (Dantzig et al., 1954).
+  2. 2.
+    Heuristics: Methods include Nearest Neighbor, Insertion Methods, and Christofides Algorithm (3/2-approximation) (Christofides, 1976).
+  3. 3.
+    Meta-heuristics: This category includes Genetic Algorithms (Holland, 1992), Simulated Annealing (Kirkpatrick et al., 1983), and Ant Colony Optimization (Dorigo & Stützle, 2004).
+D.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
+Node Components (N)
+For TSP with n locations:
+$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (8)
+    $n_{\text{route}}$: Route generation role
+    $n_{\text{dist}}$: Distance calculation role
+    $n_{\text{valid}}$: Solution validation role
+Edge Dependencies (E)
+$E=\{e_{\text{spatial}},e_{\text{sequence}},e_{\text{complete}}\}\text{ where}$ (9)
+    $e_{\text{spatial}}$: Distance constraints between locations
+    $e_{\text{sequence}}$: Visit order constraints
+    $e_{\text{complete}}$: Tour completion requirements
+Agent Assignments
+Node Agents ($A_{n}$):
+  • Route Generation Agent: Generates candidate tours
+  • Distance Calculator Agent: Computes tour lengths
+  • Solution Validator Agent: Verifies tour validity
+Edge Agents ($A_{e}$):
+  • Spatial Constraint Agent: Monitors distance feasibility
+  • Sequence Monitor Agent: Ensures valid visit order
+  • Completion Checker Agent: Verifies tour completeness
+Algorithm Selection
+Based on the size of the problem, an algorithm is selected to balance performance trade-offs and mitigate the exponential computational cost of the brute-force method.
+Validation Function
+$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$ (10)
+D.3 Experiments, From Small to Large N
+  1. 1.
+    N=5: Establish ground truth via brute force
+  2. 2.
+    N=10,20,100: Test LLM performance degradation
+  3. 3.
+    Metrics:
+    • Solution quality vs optimal
+    • Computation attempts before giving up
+    • Error recognition capability
+D.3.1 Small Campus Tour (N=5)
+Plan an optimal route for campus tour guide visiting 5 key locations:
+  • A: Admissions Office (start/end)
+  • B: Library
+  • C: Student Center
+  • D: Science Building
+  • E: Sports Complex
+Distance Matrix (minutes)
+$D=\begin{bmatrix}0&5&8&4&7\\ 5&0&6&3&8\\ 8&6&0&5&4\\ 4&3&5&0&6\\ 7&8&4&6&0\end{bmatrix}$ (11)
+Constraints
+  • Tour starts/ends at Admissions (A)
+  • Each location visited exactly once
+  • Total possible routes: (5-1)!/2 = 12
+  • Optimal solution can be verified by hand
+3.1.1. $\mathbf{W^{*}}$ Workflow Components
+Node Components (N)
+  • $n_{\text{route}}$: Generates permutations starting/ending at A.
+  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
+  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
+Edge Dependencies (E)
+  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
+  • $e_{\text{sequence}}$: Ensures visit order consistency.
+  • $e_{\text{complete}}$: Validates all 5 locations are visited.
+Agent Assignments
+  • Node Agents:
+    • Route Generation Agent (handles $n_{\text{route}}$)
+    • Distance Calculator Agent (handles $n_{\text{dist}}$)
+    • Solution Validator Agent (handles $n_{\text{valid}}$)
+  • Edge Agents:
+    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
+    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
+    • Completion Checker Agent (enforces $e_{\text{complete}}$)
+Selected Algorithm
+Brute-force.
+Validation Function
+$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$
+3.1.2. Solution Steps
+Step 1: Problem Parsing
+  • Input: 5 locations with distance matrix $D$.
+  • Initialize node/edge agents and constraints.
+Step 2: Route Generation ($n_{\text{route}}$ Agent)
+  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
+  • Example permutations:
+    • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
+    • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
+Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
+  • Compute total time for each route using $D$.
+Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
+  • Check all routes for:
+    • Start/end at A (e.g., invalid route: $A\rightarrow B\rightarrow C\rightarrow D\rightarrow E\rightarrow B$).
+    • No duplicate visits.
+Step 5: Edge Agent Validation
+  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
+  • Sequence Monitor Agent: Confirm no backtracking (e.g., $B\rightarrow D$ allowed; $D\rightarrow B$ invalid unless part of loop).
+Step 6: Apply Validation Function
+  • Assign $V=-\infty$ to invalid routes.
+[…]
+  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
+3.1.3 Solution
+Optimal tour time: 242424 minutes, achieved by three routes:
+  • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
+  • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
+  • $A\rightarrow E\rightarrow C\rightarrow B\rightarrow D\rightarrow A$
+D.3.2 Large Campus Tour (N=10)
+Plan an optimal route for a guided tour through 10 locations:
+  • Locations: A (Admissions), B (Library), C (Student Center), …, J (Sports Complex)
+  • Distance Matrix: Asymmetric travel times (minutes)
+$D=\begin{pmatrix}0&12&8&15&9&14&7&11&10&6\\ 10&0&7&14&6&16&9&13&5&8\\ 9&5&0&11&8&12&10&7&15&4\\ 14&8&12&0&10&9&13&6&11&7\\ 7&13&6&9&0&8&5&12&14&10\\ 11&9&15&8&12&0&7&10&13&5\\ 5&7&10&6&11&9&0&8&12&15\\ 8&14&4&10&7&13&6&0&9&11\\ 12&6&9&7&15&10&8&5&0&14\\ 9&10&7&13&5&11&14&8&12&0\end{pmatrix}$
+Table 7: Comparison of Planners and Their Performance Characteristics. $\mathcal{MP}$ provides validation to improve accuracy.
+Planner Best Results Algorithm Iters. Advantages Limitations
+Claude $92\rightarrow 66$ mins Nearest Neighbor 3 Efficiently implements greedy heuristic approach Makes data reading errors, compromising solution accuracy
+GPT4o $75\rightarrow 68$ mins Genetic 3 Identifies effective termination conditions Unable to implement exact algorithms like Held-Karp
+DeepSeek $60$ mins Held-Karp 1 Implements optimal algorithm correctly None observed for this problem size
+$\mathcal{MP}$ + Claude $66\rightarrow 60$ mins Ant Colony Optimization 2 Provides validation and suggests iteration increases for improvement Requires external guidance for algorithm selection and parameter tuning
+MP + GPT4o $60$ mins Ant Colony Optimization 1 Achieves optimal solution with precise execution Requires more computational resources with larger ant population and iteration count
+MP + DeepSeek $60$ mins Ant Colony Optimization 1 Combines efficient algorithm selection with optimal parameter tuning None significant for given problem
+Algorithm Selection
+Based on the size of the problem, $\mathcal{MP}$ selected the Ant Colony Optimization (Dorigo & Stützle, 2004) algorithm to achieve at least a 4x speedup.
+For $N\geq 10$, an approximate method is recommended.
+3.2.1. ACO Method
+Parameters
+  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
+  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
+The termination criteria can be modified to stop the algorithm if no meaningful improvement is observed after $k$ consecutive iterations.
+Algorithm
+1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
+2:  for 50 iterations do
+3:     for all 100 ants do
+4:        Build tour using $P_{ij}=\frac{[\tau_{ij}]^{1}[\eta_{ij}]^{2}}{\sum[\tau_{ik}]^{1}[\eta_{ik}]^{2}}$
+5:        Record tour length $L_{k}$
+6:     end for
+7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
+8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
+9:     Track best tour
+10:  end for
+3.2.2. Performance Comparison
+The goalpost is the optimal time of 606060 minutes.
+Table 7 compares six different configurations, and three of the six achieve the optimal answer.
+Although TSP is a relatively simple scheduling problem with just one actor and no parallel execution, the benefit of having $\mathcal{MP}$ to validate results is still helpful to Glaude and GPT4o.
+When asked to solve the problem without $\mathcal{MP}$, Glaude and GPT4o initially chose brute force, then switched to an approximation method without thorough deliberation (or perhaps they did, but did not output their reasoning process).
+However, DeekSeek picked Held-Karp, a computationally expensive method, even more expensive than brute force, arguing that the absolute computation time for $N=10$ is only $0.2$ seconds.
+$\mathcal{MP}$ was more deliberate, opting for brute force when $N=5$ and ACO for $N=10$.
+D.4 TSP Experiment Conclusion
+This simple task demonstrates that $\mathcal{MP}$ can be valuable for monitoring the execution process, validating the correctness of intermediate results, and suggesting more efficient algorithmic approaches.
+Appendix E Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
 The problem statement remains consistent with Section 3, with $\mathbf{W^{*}}$ generated by $\mathcal{MP}$ to enhance constraints and dependencies.
 Planning performance is compared across four configurations: DeepSeek, GPT4o, DeepSeek + $\mathcal{MP}$, and GPT4o + $\mathcal{MP}$.
-A.1 Phase 1: Network Construction
-A.1.1 Node (Role) Specifications
+E.1 Phase 1: Network Construction
+E.1.1 Node (Role) Specifications
 First, meta-planner $\mathcal{MP}$ extracts roles ($\mathcal{N}$) with their required qualifications:
@@
   • $n_{\text{supervisor}}$: capability to monitor oven
-A.1.2 Edge (Dependency) Specifications
+E.1.2 Edge (Dependency) Specifications
 Next, $\mathcal{MP}$ identifies dependencies ($\mathcal{E}$) between roles:
@@
   • $e_{\text{safety}}$: - Continuous oven supervision requirement
-A.2 Phase 2: Agent Assignments
+E.2 Phase 2: Agent Assignments
 After constructing the network structure, $\mathcal{MP}$ selects and assigns agents to monitor both the roles and dependencies.
-A.2.1 Node (Role) Agent Assignment
+E.2.1 Node (Role) Agent Assignment
 For each role, $\mathcal{MP}$ selects monitoring agents with the required capabilities:
@@
   • Resource Monitor: Manages vehicle assignments and actor schedules
-A.2.2 Edge (Dependency) Agent Assignment
+E.2.2 Edge (Dependency) Agent Assignment
 For the identified dependencies, $\mathcal{MP}$ assigns specialized monitoring agents:
@@
 Mitigation plans
-A.2.3 Common Sense Constraint Analysis (Performed by an LLM)
+E.2.3 Common Sense Constraint Analysis (Performed by an LLM)
 A common sense agent identifies the following implicit constraints that can affect Thanksgiving dinner planning. This list is generated by Claude given the problem statement.
@@
 - Critical tasks
-A.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
+E.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
 The common sense constraints identified above require different verification approaches:
@@
 $\mathcal{MP}$ can then integrate both types of information into the final workflow design.
-A.3 Agent Requirements and Assignments
+E.3 Agent Requirements and Assignments
 The $\mathcal{MP}$ requires two categories of agents. $\mathcal{MP}$ specifies their requirements in the protocol buffer format in Table 8 for the nodes and Table 9 for the edges, respectively.
@@
 During execution, agents communicate through these standardized protocols while maintaining their specialized monitoring functions.
-A.4 Monitoring Protocols and Dynamic Adjustments
+E.4 Monitoring Protocols and Dynamic Adjustments
 The workflow monitoring operates through a hierarchical protocol system that enables both routine supervision and dynamic adjustments.
@@
 The meta-planner $\mathcal{MP}$ coordinates these responses while maintaining global constraint satisfaction.
-A.5 Integrated Workflow Network
+E.5 Integrated Workflow Network
 Table 10 presents the resulting workflow network $\mathbf{W^{*}}$, which includes all nodes and edges, and their assigned agents and protocols.
@@
     • Safety Agent: Role coverage for supervision
-A.6 Agent Interaction Specifications
+E.6 Agent Interaction Specifications
 Please, see Table 11.
@@
 Alert: travel_delay()
-A.7 New Problem Statement Revised with $\mathbf{W^{*}}$
+E.7 New Problem Statement Revised with $\mathbf{W^{*}}$
 Given the $\mathbf{W^{*}}$ generated by MACI’s meta-planner $\mathcal{MP}$, the Thanksgiving Dinner Planning problem statement stated at the beginning of this section is revised as follows:
@@
     The schedule consists of three columns: time, task, and assigned person(s).
-A.8 Experiment #1: Sequential Planner
+E.8 Experiment #1: Sequential Planner
 Once after the original plan was revised by $\mathcal{MP}$ to include more specific details, clarify ambiguous explicit constraints, and define implicit constraints, the performance of the three LLMs used in the experiment improved significantly.
@@
 6:00 PM Thanksgiving dinner begins All
-A.8.1 Results: DeepSeek Wins
+E.8.1 Results: DeepSeek Wins
 Upon closer examination of the number of iterations required to produce a feasible plan, DeepSeek and Claude each required one revision (two iterations), while GPT4o required two revisions (three iterations).
@@
 6:00 PM Dinner served All
-A.8.2 Observations on Errors
+E.8.2 Observations on Errors
 Handling Long Dependencies
@@
   • Cross-iteration consistency.
-A.9 Experiment #2: Reactive Planner for Flight Delay
+E.9 Experiment #2: Reactive Planner for Flight Delay
 At 10:00 AM Eastern time, Sarah is notified that James’s flight is delayed by three hours, with a new arrival time of 4:00 PM.
@@
     The schedule consists of three columns: time, task, and assigned person(s).
-A.9.1 Results: DeepSeek Wins
+E.9.1 Results: DeepSeek Wins
 None of the LLMs cannot react appropriately to this new event without clearing their context buffers.
-As explained in Appendix A.8.2, this limitation is evident.
+As explained in Appendix E.8.2, this limitation is evident.
 The key takeaway is that for future runtime frameworks, we must ensure infrastructure support for selectively invalidating stale constraints.
@@
 6:00 PM James arrives home. Dinner served. All
-A.9.2 Observations on Errors
+E.9.2 Observations on Errors
 The initial meta-planner failed to recognize a critical opportunity in early information detection.
@@
 DeepSeek was aware of this alert in a timely manner, but Claude was not.
-A.10 Conclusion
+E.10 Conclusion
 Our concluding remark is that we may not be able to rely on LLMs alone to cover all constraints and react promptly to various alerts.
 This reinforces that the MACI architecture is on the right path to address all the aforementioned limitations of LLMs, some of which cannot be rectified.
-Appendix B Validation and Recovery Protocols
-The validation protocol implements a multi-stage process for ensuring state consistency. When any agent proposes a state change, the validation agent initiates a sequence of checks:
-$\text{validate}(s_{t}\rightarrow s_{t+1})=\begin{cases}\text{true}&\text{if all checks pass}\\ \text{false}&\text{if any check fails}\end{cases}\vspace{-.1in}$ (15)
-The protocol begins with pre-validation.
-Before a state transition starts, the validation agent queries relevant agents about preconditions.
-For a travel booking, temporal agent verifies the proposed times fit within existing schedules.
-Spatial agent confirms the physical feasibility of movements between locations.
-Role agent checks if all actors can perform their assigned functions.
-During the transition, the protocol maintains atomic operations.
-The validation agent tracks changes across all state dimensions, ensuring partial updates cannot create inconsistent states.
-If the temporal agent approves a flight time but the resource agent finds insufficient seats, the entire transition fails and rolls back.
-Post-validation examines the resulting state.
-The validation agent verifies that all constraints remain satisfied after the change.
-Common sense agent reviews the new state for practical issues that formal checks might miss.
-Strategy agent confirms the transition aligns with overall planning objectives.
-When validation fails, the protocol triggers a structured recovery process:
-$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (16)
-Recovery begins by logging the failure cause and violated constraints.
-The strategy agent then works with domain agents to generate alternative proposals that satisfy the constraints.
-This might involve relaxing non-critical constraints or exploring different approaches to meet the planning objectives.
-B.1 Operations Research Techniques in Validation Protocols
-The validation protocols described above align closely with established methods in operations research (OR). Some relevant techniques include:
-  • Constraint Programming (CP): Focuses on solving combinatorial problems by enforcing constraints, ensuring consistency across dimensions such as temporal, spatial, and resource availability (Rossi et al., 2006).
-  • Mixed-Integer Linear Programming (MILP): Optimizes decision variables subject to linear constraints and objective functions, often used in scheduling and resource allocation (Wolsey & Nemhauser, 1998).
-  • Network Flow Algorithms: Validates feasibility and optimizes flows in networks by ensuring capacity, timing, and availability constraints are satisfied (Ahuja et al., 1993).
-  • Dynamic Programming (DP): Decomposes problems into sequential subproblems, useful for validating processes like inter-terminal walking or luggage claiming time (Bellman, 1957).
-  • Monte Carlo Simulation: Simulates scenarios to validate feasibility and robustness under uncertainty (Metropolis & Ulam, 1949).
-  • Robust Optimization: Focuses on solutions that remain feasible under uncertainty, ensuring plans adapt to disruptions (Ben-Tal & Nemirovski, 2009).
-Integration in Multi-Agent Systems
-The validation agent also employs techniques from multi-agent systems, such as:
-  • Blackboard Systems: A shared workspace for collaborative validation by different agents, ensuring global consistency (England & Engelmore, 1987).
-  • Consensus Protocols: Used for distributed validation, where agents negotiate to ensure all constraints are met (Ren & Beard, 2005).
-By combining these OR techniques with agent-based systems, the validation protocol ensures comprehensive and adaptive checks for workflow consistency.
-Future work can explore integrating heuristic methods, such as genetic algorithms or simulated annealing, to further enhance recovery processes.
-Appendix C $\mathsf{MACI}$ Additional Design Considerations
-C.1 Cross-Domain Generalization
-While the state space dimensions—Who, Where, When, What, and Why—are broad enough to cover diverse domains, additional customization may be required for unique applications.
-This section examines how MACI generalizes across domains like financial planning, healthcare logistics, and supply chain optimization.
-The travel planning example is illustrative, emphasizing how MACI dynamically adapts state spaces and agents to domain-specific requirements.
-C.2 Dynamic Agent Registration and Evolution
-This section explores how agents are dynamically developed, trained, and validated for new tasks.
-It discusses mechanisms for evaluating new agents and integrating them into the repository without retraining the entire system, ensuring scalability and adaptability.
-C.3 Scalability and Resource Efficiency
-As the number of agents and task complexity grows, MACI employs strategies to manage communication overhead and optimize agent interactions.
-This section details techniques for clustering agents and hierarchical coordination to maintain scalability.
-C.4 Empirical Evaluation Across Domains
-To demonstrate MACI’s adaptability, this section presents empirical results from applying the framework to multiple domains.
-Examples include financial portfolio management, urban traffic planning, and hospital resource allocation, highlighting MACI’s advantages over state-of-the-art systems.
-C.5 Challenges and Future Directions
-While MACI addresses many limitations of LLM-based planning, challenges remain in real-time coordination, implicit knowledge integration, and robust recovery mechanisms.
-This section proposes future research directions to enhance MACI’s performance and applicability to novel tasks.
-Appendix D Additional Tables and Figures
-D.1 Notation
-Table 20 presents all symbols used throughout this paper.
-Table 20: Symbol Definitions
-Symbol Definition Symbol Definition
-Basic Sets
-$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
-$\mathcal{C_{E}}$ Explicit constraints $\mathcal{C_{I}}$ Implicit constraints
-$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
-Workflow Components
-$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
-$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
-$n_{i}$ Individual role $e_{ij}$ Role dependency
-Agent Functions
-$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
-$V(\cdot)$ Validation function $\text{dist}(\cdot)$ Capability distance
-$\text{map}_{\text{role}}$ Role extraction $\text{map}_{\text{edge}}$ Dependency extraction
-Optimization
-$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
-$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
-D.2 Example Common Agents
-  1. 1.
-    Role Manager Agent (Who): Tracks actors, their roles, and their associated constraints, ensuring that all role-based requirements are satisfied.
-  2. 2.
-    Spatial Agent (Where): Manages geographic and location-based constraints, verifying transitions between physical or virtual locations.
-  3. 3.
-    Temporal Agent (When): Handles scheduling, timing, and deadlines, ensuring alignment with temporal constraints.
-  4. 4.
-    Resource Agent (What): Tracks real-world resources such as tools, vehicles, or financial instruments, managing capacity, availability, and associated costs.
-  5. 5.
-    Reasoning and Explanation Agent (Why): Maintains the rationale behind decisions, dependencies, and alternative plans, enabling consistent alignment with objectives and providing explanations for outcomes.
-  6. 6.
-    Common Sense Agent: Identifies implicit constraints, integrates practical knowledge, and ensures plans align with real-world considerations.
-  7. 7.
-    Constraint Validation Agent: Ensures that all constraints are satisfied and that proposed plans remain feasible.
-  8. 8.
-    Plan Evaluation Agent: Assesses the effectiveness of plans against predefined metrics and objectives.
-  9. 9.
-    What-If Testing Agent: Evaluates plan robustness by simulating alternative scenarios and analyzing their impact.
-  10. 10.
-    Compliance and Safety Agent: Monitors adherence to safety standards, ethical principles, and regulatory frameworks.
-Appendix E Traveling Salesman Problem Experiment
-E.1 General Problem Specification
-The TSP requires finding the shortest possible route visiting N locations exactly once, returning to the start:
-  • Inputs: N locations, distance matrix D[N][N]
-  • Output: Optimal tour T minimizing total distance
-  • Constraints: Each location visited once, return to start
-Computational Complexity - Brute Force
-For N locations:
-  • Number of possible tours = (N-1)!/2
-  • Time complexity = O(N!)
-  • Space complexity = O(N2)
-Solution Methods
-  1. 1.
-    Exact Methods: Representative methods are Branch and Bound (Land & Doig, 1960), Dynamic Programming (Bellman, 1962), and Integer Linear Programming (Dantzig et al., 1954).
-  2. 2.
-    Heuristics: Methods include Nearest Neighbor, Insertion Methods, and Christofides Algorithm (3/2-approximation) (Christofides, 1976).
-  3. 3.
-    Meta-heuristics: This category includes Genetic Algorithms (Holland, 1992), Simulated Annealing (Kirkpatrick et al., 1983), and Ant Colony Optimization (Dorigo & Stützle, 2004).
-E.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
-Node Components (N)
-For TSP with n locations:
-$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (17)
-    $n_{\text{route}}$: Route generation role
-    $n_{\text{dist}}$: Distance calculation role
-    $n_{\text{valid}}$: Solution validation role
-Edge Dependencies (E)
-$E=\{e_{\text{spatial}},e_{\text{sequence}},e_{\text{complete}}\}\text{ where}$ (18)
-    $e_{\text{spatial}}$: Distance constraints between locations
-    $e_{\text{sequence}}$: Visit order constraints
-    $e_{\text{complete}}$: Tour completion requirements
-Agent Assignments
-Node Agents ($A_{n}$):
-  • Route Generation Agent: Generates candidate tours
-  • Distance Calculator Agent: Computes tour lengths
-  • Solution Validator Agent: Verifies tour validity
-Edge Agents ($A_{e}$):
-  • Spatial Constraint Agent: Monitors distance feasibility
-  • Sequence Monitor Agent: Ensures valid visit order
-  • Completion Checker Agent: Verifies tour completeness
-Algorithm Selection
-Based on the size of the problem, an algorithm is selected to balance performance trade-offs and mitigate the exponential computational cost of the brute-force method.
-Validation Function
-$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$ (19)
-E.3 Experiments, From Small to Large N
-  1. 1.
-    N=5: Establish ground truth via brute force
-  2. 2.
-    N=10,20,100: Test LLM performance degradation
-  3. 3.
-    Metrics:
-    • Solution quality vs optimal
-    • Computation attempts before giving up
-    • Error recognition capability
-E.3.1 Small Campus Tour (N=5)
-Plan an optimal route for campus tour guide visiting 5 key locations:
-  • A: Admissions Office (start/end)
-  • B: Library
-  • C: Student Center
-  • D: Science Building
-  • E: Sports Complex
-Distance Matrix (minutes)
-$D=\begin{bmatrix}0&5&8&4&7\\ 5&0&6&3&8\\ 8&6&0&5&4\\ 4&3&5&0&6\\ 7&8&4&6&0\end{bmatrix}$ (20)
-Constraints
-  • Tour starts/ends at Admissions (A)
-  • Each location visited exactly once
-  • Total possible routes: (5-1)!/2 = 12
-  • Optimal solution can be verified by hand
-3.1.1. $\mathbf{W^{*}}$ Workflow Components
-Node Components (N)
-  • $n_{\text{route}}$: Generates permutations starting/ending at A.
-  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
-  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
-Edge Dependencies (E)
-  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
-  • $e_{\text{sequence}}$: Ensures visit order consistency.
-  • $e_{\text{complete}}$: Validates all 5 locations are visited.
-Agent Assignments
-  • Node Agents:
-    • Route Generation Agent (handles $n_{\text{route}}$)
-    • Distance Calculator Agent (handles $n_{\text{dist}}$)
-    • Solution Validator Agent (handles $n_{\text{valid}}$)
-  • Edge Agents:
-    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
-    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
-    • Completion Checker Agent (enforces $e_{\text{complete}}$)
-Selected Algorithm
-Brute-force.
-Validation Function
-$V(W,M)=\begin{cases}-\infty&\text{if constraints violated}\\ -\text{tour\_length}&\text{if tour valid}\end{cases}$
-3.1.2. Solution Steps
-Step 1: Problem Parsing
-  • Input: 5 locations with distance matrix $D$.
-  • Initialize node/edge agents and constraints.
-Step 2: Route Generation ($n_{\text{route}}$ Agent)
-  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
-  • Example permutations:
-    • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
-    • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
-Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
-  • Compute total time for each route using $D$.
-Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
-  • Check all routes for:
-    • Start/end at A (e.g., invalid route: $A\rightarrow B\rightarrow C\rightarrow D\rightarrow E\rightarrow B$).
-    • No duplicate visits.
-Step 5: Edge Agent Validation
-  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
-  • Sequence Monitor Agent: Confirm no backtracking (e.g., $B\rightarrow D$ allowed; $D\rightarrow B$ invalid unless part of loop).
-Step 6: Apply Validation Function
-  • Assign $V=-\infty$ to invalid routes.
-[…]
-  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
-3.1.3 Solution
-Optimal tour time: 242424 minutes, achieved by three routes:
-  • $A\rightarrow D\rightarrow B\rightarrow C\rightarrow E\rightarrow A$
-  • $A\rightarrow B\rightarrow D\rightarrow C\rightarrow E\rightarrow A$
-  • $A\rightarrow E\rightarrow C\rightarrow B\rightarrow D\rightarrow A$
-E.3.2 Large Campus Tour (N=10)
-Plan an optimal route for a guided tour through 10 locations:
-  • Locations: A (Admissions), B (Library), C (Student Center), …, J (Sports Complex)
-  • Distance Matrix: Asymmetric travel times (minutes)
-$D=\begin{pmatrix}0&12&8&15&9&14&7&11&10&6\\ 10&0&7&14&6&16&9&13&5&8\\ 9&5&0&11&8&12&10&7&15&4\\ 14&8&12&0&10&9&13&6&11&7\\ 7&13&6&9&0&8&5&12&14&10\\ 11&9&15&8&12&0&7&10&13&5\\ 5&7&10&6&11&9&0&8&12&15\\ 8&14&4&10&7&13&6&0&9&11\\ 12&6&9&7&15&10&8&5&0&14\\ 9&10&7&13&5&11&14&8&12&0\end{pmatrix}$
-Table 21: Comparison of Planners and Their Performance Characteristics. $\mathcal{MP}$ provides validation to improve accuracy.
-Planner Best Results Algorithm Iters. Advantages Limitations
-Claude $92\rightarrow 66$ mins Nearest Neighbor 3 Efficiently implements greedy heuristic approach Makes data reading errors, compromising solution accuracy
-GPT4o $75\rightarrow 68$ mins Genetic 3 Identifies effective termination conditions Unable to implement exact algorithms like Held-Karp
-DeepSeek $60$ mins Held-Karp 1 Implements optimal algorithm correctly None observed for this problem size
-$\mathcal{MP}$ + Claude $66\rightarrow 60$ mins Ant Colony Optimization 2 Provides validation and suggests iteration increases for improvement Requires external guidance for algorithm selection and parameter tuning
-MP + GPT4o $60$ mins Ant Colony Optimization 1 Achieves optimal solution with precise execution Requires more computational resources with larger ant population and iteration count
-MP + DeepSeek $60$ mins Ant Colony Optimization 1 Combines efficient algorithm selection with optimal parameter tuning None significant for given problem
-Algorithm Selection
-Based on the size of the problem, $\mathcal{MP}$ selected the Ant Colony Optimization (Dorigo & Stützle, 2004) algorithm to achieve at least a 4x speedup.
-For $N\geq 10$, an approximate method is recommended.
-3.2.1. ACO Method
-Parameters
-  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
-  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
-The termination criteria can be modified to stop the algorithm if no meaningful improvement is observed after $k$ consecutive iterations.
-Algorithm
-1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
-2:  for 50 iterations do
-3:     for all 100 ants do
-4:        Build tour using $P_{ij}=\frac{[\tau_{ij}]^{1}[\eta_{ij}]^{2}}{\sum[\tau_{ik}]^{1}[\eta_{ik}]^{2}}$
-5:        Record tour length $L_{k}$
-6:     end for
-7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
-8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
-9:     Track best tour
-10:  end for
-3.2.2. Performance Comparison
-The goalpost is the optimal time of 606060 minutes.
-Table 21 compares six different configurations, and three of the six achieve the optimal answer.
-Although TSP is a relatively simple scheduling problem with just one actor and no parallel execution, the benefit of having $\mathcal{MP}$ to validate results is still helpful to Glaude and GPT4o.
-When asked to solve the problem without $\mathcal{MP}$, Glaude and GPT4o initially chose brute force, then switched to an approximation method without thorough deliberation (or perhaps they did, but did not output their reasoning process).
-However, DeekSeek picked Held-Karp, a computationally expensive method, even more expensive than brute force, arguing that the absolute computation time for $N=10$ is only $0.2$ seconds.
-$\mathcal{MP}$ was more deliberate, opting for brute force when $N=5$ and ACO for $N=10$.
-E.4 TSP Experiment Conclusion
-This simple task demonstrates that $\mathcal{MP}$ can be valuable for monitoring the execution process, validating the correctness of intermediate results, and suggesting more efficient algorithmic approaches.